### v2

# SA-Person: Text-Based Person Retrieval with Scene-aware Re-ranking

Abstract
Text-based person retrieval aims to identify a target individual from a gallery of images based on a natural language description. It presents a significant challenge due to the complexity of real-world scenes and the ambiguity of appearance-related descriptions. Existing methods primarily emphasize appearance-based cross-modal retrieval, often neglecting the contextual information embedded within the scene, which can offer valuable complementary insights for retrieval. To address this, we introduce ScenePerson-13W, a large-scale dataset featuring over 100,000 scenes with rich annotations covering both pedestrian appearance and environmental cues. Based on this, we propose SA-Person, a two-stage retrieval framework. In the first stage, it performs discriminative appearance grounding by aligning textual cues with pedestrian-specific regions. In the second stage, it introduces SceneRanker, a training-free, scene-aware re-ranking method leveraging multimodal large language models to jointly reason over pedestrian appearance and the global scene context. Experiments on ScenePerson-13W validate the effectiveness of our framework in challenging scene-level retrieval scenarios. The code and dataset will be made publicly available.
00footnotetext: * These authors contributed equally to this work.00footnotetext: †Corresponding author.
1 Introduction
Person retrieval galiyawala2021person aims to identify a target individual within the image galleries based on a query, playing a vital role in applications like video surveillance and human-centric artificial intelligent systems li2018richly ; galiyawala2021person ; li2023dcel . Early efforts focus on image-based paradigms zhang2016learning ; liu2018pose ; sarfraz2018pose ; xiao2017joint ; he2018end ; yan2021anchor ; cao2022pstr ; yan2022exploring using cropped images as queries, and then shift toward text-based retrieval li2017person ; yu2019cross ; liu2023survey due to the accessibility of textual descriptions. Advances in vision-language pretraining (VLP) models, such as CLIP radford2021learning , also accelerate its development yang2023towards ; li2024adaptive ; zuo2024ufinebench ; luo2025graph ; liu2025dm ; jiang2023cross ; bai2023rasa ; cao2024empirical ; ergasti2024mars ; qin2024noisy . However, text-based methods rely solely on cropped images and textual descriptions of appearance, limiting their effectiveness in real-world scenarios. Recent methods zhang2023text ; su2024maca ; yan2025fusionsegreid extend retrieval to full-scene images but still focus narrowly on appearance, neglecting critical environmental scene cues. This limitation poses challenges in complex scenarios with similar-looking individuals, or the same person under varying contexts. Incorporating contextual scene semantics alongside individual appearance is essential for accurate and robust person retrieval.
This work focuses on scene-aware text-based person retrieval. The first challenge is the lack of a dedicated dataset for this task. To this end, we introduce ScenePerson-13W, a novel dataset designed for real-world scene-aware person retrieval. It includes over $131,438$ images from $12,562$ YouTube video sequences, featuring diverse individuals across varying crowd densities, camera angles, and environments. Unlike existing datasets that focus on cropped images or weakly grounded captions, ScenePerson-13W provides rich image-text pairs with detailed descriptions encompassing both visual appearance and contextual scene information, which promotes text-to-person alignment.
Building on ScenePerson-13W, we propose SA-Person, a two-stage framework for scene-aware person retrieval. In the first stage, appearance-related information is extracted from textual descriptions to identify candidate pedestrians via cross-modal matching, reducing the retrieval space with discriminative appearance features. In the second stage, we introduce SceneRanker, a training-free, scene-aware re-ranking module powered by MLLM. By employing structured multimodal prompts, SceneRanker enables holistic reasoning over appearance and global scene context, allowing effective disambiguation of visually similar individuals and significantly improving retrieval performance.
In summary, our main contributions are as follows:
  • We introduce ScenePerson-13W, a new large-scale dataset for scene-aware text-based person retrieval, with compositional descriptions that include both pedestrian appearance and global scene context, alleviating the limitations of prior datasets.
  • We present SA-Person, a two-stage retrieval framework that combines discriminative appearance grounding with holistic scene understanding, enabling robust retrieval from large-scale full-scene image databases.
  • We propose SceneRanker, the first re-ranking strategy for text-based person retrieval powered by MLLM, which follows a training-free pattern and incorporates visual grounding and contextual reasoning to improve retrieval accuracy in complex scenes.
  • Extensive experiments on ScenePerson-13W demonstrate the effectiveness and generalizability of our proposed SA-Person across various retrieval settings and model backbones.
Figure 1: Overview of the ScenePerson-13W construction pipeline. The construction pipeline involves scene segmentation, pedestrian detection and tracking, completeness filtering, image deduplication, and textual description generation. For each retained pedestrian, a description is generated based on the full image with a highlighted target, capturing their appearance, spatial location, and relationships with surrounding elements.
Table 1: Statistics of person retrieval datasets. SCENEPERSON-13W shows a substantial advancement over existing person retrieval datasets by combining fine-grained appearance details with rich global scene contexts.
Dataset Query Modality Gallery Modality #Description #Image
Market-1501 zheng2015scalable Image (Cropped) Image (Cropped) 32,668
CUHK03 li2014deepreid Image (Cropped) Image (Cropped) 14,097
CUHK-SYSU xiao2017joint Image (Cropped) Image (Full) 18,184
PRW zheng2017person Image (Cropped) Image (Full) 11,816
CUHK-PEDES li2017person Text (Appearance) Image (Cropped) 80,412 40,206
ICFG-PEDES ding2021semantically Text (Appearance) Image (Cropped) 54,522 54,522
CUHK-SYSU-TBPS zhang2023text Text (Appearance) Image (Full) 35,960 18,184
PRW-TBPS zhang2023text Text (Appearance) Image (Full) 19,009 11,246
ScenePerson-13W Text (Appearance+scene) Image (Full) 131,438 131,438
2 ScenePerson-13W Dataset
The ScenePerson-13W dataset is proposed to support scene-aware person retrieval by incorporating rich scene context alongside detailed pedestrian appearance in textual descriptions. It contains $131,438$ pedestrian instances across $112,872$ diverse scenes, with each instance accompanied by a textual description, a full-scene image, and a bounding box highlighting the pedestrian’s location. The dataset is split into a training set of $121,438$ images and a testing set of $10,000$ images, which preserves the complex scene structures and identity distributions while reducing redundancy and ensures comprehensive evaluation across diverse scenarios. The construction process of the dataset illustrated in Figure 1, involves five key steps: scene segmentation, pedestrian detection and tracking, completeness filtering, deduplication in image collection, and textual description generation. More examples from the dataset are presented in Appendix A.3.
Image Collection. The image data in ScenePerson-13W is constructed from the QVHighlights video dataset lei2107qvhighlights through a multi-stage processing pipeline designed to ensure data quality and diversity. First, videos are segmented into distinct scenes, each representing a unique environmental context. This segmentation enables the dataset to capture a wide range of pedestrian-environment combinations, supporting comprehensive scene-aware analysis. Second, the pedestrian detection and tracking are performed using YOLOv11 khanam2024yolov11 . For each detected pedestrian instance, the frame with the highest confidence score is selected, and the corresponding full scene image with cropped pedestrian region are retained. Third, to enhance data quality, incomplete or noisy detections–such as those capturing only partial body parts–are filtered out using MediaPipe keypoint detection mediapipe_pose . Only images containing a head and at least one shoulder are preserved, ensuring the availability of essential visual features. Finally, a deduplication process is applied to reduce redundancy. Using CLIP radford2021learning , visual similarity between images is assessed. Instances that are highly similar in both the cropped pedestrian region and the full scene are discarded, promoting the uniqueness and diversity of the resulting dataset. Finally, we collect $131,438$ pedestrian instance, accompanied with $131,438$ corresponding scene images.
Description Generation. For each of the $131,438$ retained pedestrian instances, a corresponding textual description is generated using InternVL2.5-8B chen2024expanding . Specifically, the model is prompted with the full-scene image in which the target pedestrian is highlighted using a red bounding box. The prompt is structured as follows:
Generate a description of the individual within the red bounding box and the connection with the surroundings, in a style similar to the template: {template}. Ensure that the description is concise within $50$ words and accurate. Do not include text details or labels, and strictly avoid any punctuation other than commas and periods.
The prompt is designed to elicit detailed descriptions by guiding the model to focus on three key aspects: the individual’s appearance, location within the scene, and relationships with surrounding elements, following a structured template.
Compared to existing person retrieval datasets, ScenePerson-13W offers a significant advancement by integrating fine-grained appearance details with rich global scene context across diverse scenarios, as shown in Table  1.
Figure 2: Overview of the proposed SA-Person framework. The first stage, appearance-based person retrieval, aligns pedestrian-specific regions with appearance-related descriptions for initial retrieval. The second stage, scene-aware reranking, re-ranks the top-$K$ candidates using scene-aided text and the full-scene image.
3 Method
To tackle the challenges of holistic understanding in both pedestrian appearance and scene context, we propose SA-Person, a two-stage framework, as shown in Figure 2. The first stage aligns pedestrian-specific regions with appearance-related descriptions to retrieve candidates. The second stage refines the initial candidates by incorporating global scene context.
3.1 Appearance-based Retrieval
Appearance Embedding Training. The first stage focuses on retrieving a shortlist of candidate individuals from the gallery set based solely on their visual appearance, given the appearance-related content of the textual query. For a gallery of full-scene images $\mathcal{G}=\{I_{1},I_{2},\dots,I_{N}\}$ and a textual query $T$ describing the target person, we begin by extracting the appearance-related content from $T$. Specifically, we parse each query description in the dataset and extract the segments of the description that describes the target person’s visual traits, such as clothing, hairstyle, or accessories (e.g., “A man wearing a red jacket and blue jeans is laughing”), denoted as ${T}_{\text{app}}$. This ensures that the retrieval process prioritizes appearance-focused matching, excluding contextual or environmental cues at this stage.
To process the visual input $\mathcal{G}$, we employ an off-the-shelf pedestrian detector to identify and crop all individuals within each scene image $I_{i}$. This results in a set of bounding boxes $\mathcal{B}_{i}=\{B_{i}^{1},B_{i}^{2},\dots,B_{i}^{m_{i}}\}$, where $B_{i}^{j}=(x_{j},y_{j},w_{j},h_{j})$ represents the coordinates of the $j$-th detected person, ${m_{i}}$ represents the number of person detected in $I_{i}$. Each cropped region $B_{i}^{j}$ is treated as an independent person image for feature extraction.
We employ a dual-encoder architecture, initialized with a pre-trained vision-language model, to encode the cropped person images and appearance-related textual descriptions. For each cropped person image $b_{ij}$ and its corresponding appearance-related textual description ${T}_{\text{app}}$, the visual and textual feature is computed as:
$f_{ij}^{v}=\text{Enc}_{\text{v}}(b_{ij}),\quad f^{t}=\text{Enc}_{\text{t}}({T}_{\text{app}}),$ (1)
where $\text{Enc}_{\text{v}}$ and $\text{Enc}_{\text{t}}$ represent visual encoder and textual encoder, respectively, $f_{ij}^{v}\in\mathbb{R}^{d}$ represents the visual feature representation of the $j$-th person in image $I_{i}$ and $f^{t}\in\mathbb{R}^{d}$ is the textual feature representation aligned with the visual feature space.
To enhance cross-modal alignment, we fine-tune the dual-encoder on these appearance-related image-text pairs from ScenePerson-13W, following the implicit relation reasoning proposed in jiang2023cross . The loss is formally expressed as:
$\mathcal{L}_{\text{app}}=\mathbb{E}_{(b_{ij},\mathcal{T}_{\text{app}},y_{ij})\sim\mathcal{D}_{\text{app}}}\left[\ell_{\text{IRR}}(f_{ij}^{v},f^{t},y_{ij};\Theta)+\ell_{\text{SDM}}(f_{ij}^{v},f^{t},y_{ij};\Theta)+\ell_{\text{ID}}(f_{ij}^{v},f^{t},y_{ij};\Theta)\right],$ (2)
where $\mathcal{D}_{\text{app}}$ is the set of appearance-related image-text pairs with identity labels $y_{ij}$, $\ell_{\text{IRR}}$, $\ell_{\text{SDM}}$, and $\ell_{\text{ID}}$ are the respective loss components as defined in  jiang2023cross , and $\Theta$ denotes the parameters of the image and text encoders. The expectation is computed over the training set, ensuring discriminative and robust feature representations.
Retrieval. During inference, we first apply a YOLO-based person detector to a gallery of full-scene images $\mathcal{G}=\{\mathbf{I}_{j}\}_{j=1}^{M}$, extracting all detected person crops across the gallery. This yields a set of $K$ cropped person images $\mathcal{C}=\{\mathbf{C}_{k}\}_{k=1}^{K}$, where each $\mathbf{C}_{k}$ originates from a full-scene image $\mathbf{I}_{j}\in\mathcal{G}$. For each crop $\mathbf{C}_{k}$, we also retain its bounding box $\mathbf{B}_{k}$ and its source image identifier $j=s(k)$, indicating that $\mathbf{C}_{k}$ was cropped from $\mathbf{I}_{s(k)}$.
Each cropped image $C_{k}$ is encoded into $f_{k}^{v}$ using the fine-tuned image encoder and the appearance-related textual query $\mathcal{T}_{\text{app}}$ is encoded into $f^{t}$, as defined previously in Eq.(1). The similarity score between the textual feature and each person image is computed as:
$S_{k}=s(\mathbf{f}^{t},\mathbf{f}_{k}^{v})=\frac{\mathbf{{f}^{t}}^{\top}\mathbf{f}_{k}^{v}}{\|\mathbf{f}^{t}\|\|\mathbf{f}_{k}^{v}\|}.$ (3)
The top $K$ candidates with the highest similarity scores are selected:
$\mathcal{C}_{\text{top}}=\left\{(\mathbf{C}_{k},\mathbf{B}_{k},\mathbf{I}_{s(k)})\mid k\in\text{argmax}_{K}S_{k}\right\}.$ (4)
$\mathcal{C}_{\text{top}}$ will be passed to the subquent re-ranking stage for further improvement based on scene understanding.
3.2 Scene-aware Re-ranking
While the appearance-based retrieval stage provides a strong initial filter by matching at the region-level focused on appearance, it lacks the ability to incorporate broader contextual cues, such as spatial configurations, interactions, or referential expressions. To complement this, we introduce a scene-aware ranking module based on InternVL 2.5-8B chen2024expanding , an advanced multimodal large language model (MLLM) capable of joint visual-linguistic reasoning over complex scenes and full descriptions, without any task-specific fine-tuning.
Candidate-to-Scene Mapping. After obtaining a ranked list of cropped person candidates from the coarse retrieval stage, we map each crop back to its corresponding full-scene image. Specifically, we link each candidate crop $\mathbf{C}_{m}$ to the source image $\mathbf{I}_{s(m)}\in\mathcal{G}$ it was extracted from, along with its corresponding bounding box by the source index $s(m)$. This mapping step bridges the pedestrian-specific region and its broader scene context.
The associated bounding box $\mathbf{B}_{m}$ specifies the location of the person within $\mathbf{I}_{m}$. It tends to serve as a visual prompt that anchors the model’s attention on each candidate individual while enabling contextual reasoning about their surrounding scene, such as spatial configurations, co-occurring objects, and relational cues. This scene-aware grounding allows the model to evaluate not only the appearance of the person but also how well their broader context aligns with the textual description.
Scene-aware Multimodal Prompt. To guide SceneRanker to perform reranking, we construct a unified prompt that integrates both the annotated visual inputs and an instructional textual query. The MLLM input is a structured multimodal prompt combining the $K$ full-scene images $\{I_{m}\}_{m=1}^{K}$, their corresponding bounding box annotations $\{\mathbf{B}_{m}\}_{m=1}^{K}$, and the complete textual query $T$. The visual prompt is formatted as a markdown sequence, where each full-scene image is paired with its bounding box annotation:
$\mathbf{P}_{v}=\text{Concat}\left(\{\text{Image-}m:\langle\text{image}\rangle\langle\text{box}\rangle\mathbf{B}_{m}\langle/\text{box}\rangle\}_{m=1}^{K}\right),$
[…]
Image-1: <image><box>$B_{1}$</box> Image-2: <image><box>$B_{2}$</box> Image-$K$: <image><box>$B_{K}$</box> For the text description: ‘$T$’, order the images based on how accurately they reflect overall context in the text, starting with the most faithful match. Please output only a Python list format like this: [4, 2, 8, 1, 5, 3, 6, 9, 10, 7, $\cdots$, $K$].
Re-ranking. The SceneRanker processes the multimodal input ($(\{\mathbf{I}_{m},\mathbf{B}_{m}\}_{m=1}^{K},\mathbf{P})$ to produce a ranked list of indices:
$\mathbf{R}=\text{SceneRanker}(\{\mathbf{I}_{m},\mathbf{B}_{m}\}_{m=1}^{K},\mathbf{P}),$
where $\mathbf{R}=[m_{1},m_{2},\ldots,m_{10}]$ represents the reordered indices of the $K$ candidates, with $m_{1}$ being the index of most semantically aligned with $T$.
This stage enables the model to reason over contextual cues such as spatial layout, actions, and inter-person relations, offering complementary insights that are absent in appearance-based retrieval. The reranking output serves as the final retrieval results of our framework, integrating both localized appearance matching and global scene understanding.
4 Experiments
4.1 Experimental Setup
Dataset and Evaluation Metrics. All experiments are conducted on the proposed ScenePerson-13W dataset, which is divided into $121,438$ image-text pairs for training and $10,000$ image-text pairs for testing. Following prior retrieval works bai2023rasa ; cao2024empirical , we evaluate overall retrieval performance using Rank@k (R@k) for k=1, 5, and 10, and mean Average Precision (mAP). R@k measures the proportion of queries where the target individual is correctly identified within the top k candidates, while mAP quantifies the ranking quality across all queries.
Implementation Details. For the appearance-based retrieval stage, we adopt YOLOv5wu2021application for pedestrian detection and MediaPipe mediapipe_pose for filtering to ensure valid person instances. The dual-encoder model is fine-tuned following the original IRRA jiang2023cross implementation, initialized with CLIP-ViT-B/16 radford2021learning for images and CLIP text Transformer radford2021learning for text. The model is trained for $60$ epochs on the ScenePerson-13W dataset using the Adam optimizer with an initial learning rate of $10^{-5}$ and a cosine learning rate decay schedule. The image resolution of input crops is resized to $384\times 128$ pixels. During inference, the size of candidate set $K$ selected for further processing is set to $10$. For the SceneRanker, we use InternVL 2.5-8B chen2024expanding without task-specific fine-tuning. The generation is performed with a maximum output length of $128$ tokens and deterministic decoding. The maximum number of patches that an input image is divided into during preprocessing is set to $10$ to balance computational efficiency and detailed visual analysis. All other generation hyperparameters retain their default values as defined by the Hugging Face Transformers wolf2019huggingface library. All experiments are conducted on four RTX A40 GPUs with 48 GB of memory each.
4.2 Comparison with State-of-the-art Methods
We compare the proposed SA-Person method with a range of state-of-the-art retrieval baselines, grouped by their retrieval strategies and visual grounding scope. Region-level methods solely focus on cropped pedestrian images, emphasizing appearance cues but lacking broader scene context. Global-level methods operate on the entire scene with full textual descriptions, capturing contextual information but often sacrificing accurate grounding of pedestrian appearance. In contrast, SA-Person adopts a region-to-global retrieval paradigm, initially matching pedestrian-specific regions to ensure discriminative grounding, followed by scene-aware reranking to incorporate semantic information from the broader environment. As shown in Table 2, SA-Person consistently outperforms all baselines by a large margin. For region-level methods, it yields a significant boost of 15.41% at R@1 and 10.43% at mAP. Against the strongest region-level model, IRRA (ViT-L/14), SA-Person still improves R@1 by 14.97% and mAP by 10.15%. For global-level methods, SA-Person exceeds CLIP (ViT-L/14) by 14.83% at R@1 and 9.53 at mAP, and outperforms the best-performing LongCLIP (ViT-L/14) by 12.59% at R@1 and 7.81% at mAP. These results collectively highlight the superiority of our region-to-global cascade retrieval strategy, which integrates region-level discrimination with scene-level context and enhance both localization and semantic reasoning for more robust text-to-person retrieval.
Table 2: Comparison of the proposed SA-Person with state-of-the-art methods.
Method Focus Backbone R@1 (%) R@5 (%) R@10 (%) mAP
Region-level:
RaSa bai2023rasa Appearance ViT-B/16 62.51 83.83 88.09 72.05
IRRA jiang2023cross Appearance ViT-B/16 60.57 83.35 87.86 70.65
IRRA jiang2023cross Appearance ViT-L/14 63.37 84.54 88.67 72.71
TBPS-CLIP cao2024empirical Appearance ViT-B/16 50.29 75.24 81.58 61.54
RDE qin2024noisy Appearance ViTB/16 61.26 83.50 88.23 71.17
MARS ergasti2024mars Appearance ViTB/16 62.93 84.26 88.65 72.43
Global-level:
CLIP radford2021learning Scene-aided ViT-B/16 56.37 80.73 87.16 67.22
CLIP radford2021learning Scene-aided ViT-L/14 63.51 85.53 90.28 73.33
LongCLIP zhang2024long Scene-aided ViT-B/16 58.90 82.41 88.23 69.42
LongCLIP zhang2024long Scene-aided ViT-L/14 65.75 86.22 90.95 75.05
Region-to-Global Cascade:
SA-Person (Ours) Scene-aided ViT-B/16 77.42 86.61 87.86 81.96
SA-Person (Ours) Scene-aided ViT-L/14 78.34 87.54 88.67 82.86
Table 3: Performance comparison of text-based person retrieval with and without the proposed SceneRanker. Red numbers indicate the absolute improvement achieved by our SceneRanker.
     Methods      R@1      R@5      R@10      mAP
     TBPS-CLIP cao2024empirical      50.29      75.24      81.58      61.54
      +SceneRanker      $72.25_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+21.96}}$      $80.23_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+4.99}}$      $81.58_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+0.00}}$      $76.54_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+15.00}}$
     IRRA(ViT-B/16) jiang2023cross      60.57      83.35      87.86      70.65
      +SceneRanker      $77.42_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+16.85}}$      $86.61_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+3.26}}$      $87.86_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+0.00}}$      $81.96_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+11.31}}$
     IRRA(ViT-L/14) jiang2023cross      63.37      84.54      88.67      72.71
      +SceneRanker      $\textbf{78.34}_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+14.97}}$      $\textbf{87.54}_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+3.00}}$      $\textbf{88.67}_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+0.00}}$      $\textbf{82.86}_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+10.15}}$
     RDE qin2024noisy      61.26      83.50      88.23      71.17
      +SceneRanker      $77.58_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+16.32}}$      $86.84_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+3.34}}$      $88.23_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+0.00}}$      $82.18_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+11.01}}$
     RaSa bai2023rasa      62.51      83.83      88.09      72.05
      +SceneRanker      $77.84_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+15.33}}$      $86.83_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+3.00}}$      $88.09_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+0.00}}$      $82.27_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+10.22}}$
     MARS ergasti2024mars      62.93      84.26      88.65      72.43
      +SceneRanker      $78.32_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+15.39}}$      $87.39_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+3.13}}$      $88.65_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+0.00}}$      $82.77_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+10.34}}$
4.3 Compatibility of the Proposed SceneRanker
The proposed SceneRanker is highly compatible and can be seamlessly adapted to various text-based person retrieval methods. Specifically, these methods can function as the first stage of appearance-based retrieval in the SA-Person framework. We evaluate the compatibility of SceneRanker in Table 3, where it consistently demonstrates performance improvements. Across all baselines, SceneRanker delivers over a 15% improvement in R@1 and more than a 10% increase in mAP. Notably, it achieves significant gains for TBPS-CLIP, with R@1 increasing by 21.93% (from 50.32% to 72.25%). These consistent enhancements highlight SceneRanker as a flexible and effective component that integrates seamlessly with diverse text-based person retrieval methods. By leveraging complementary contextual cues beyond standalone pedestrian appearance, SceneRanker refines candidate selection and significantly improves retrieval results.
4.4 Ablation Studies
We conduct ablation studies to evaluate the effectiveness of the proposed SceneRanker, focusing on (1) the prompt design strategy, (2) the choice of multimodal large language models (MLLMs), and (3) the candidate size in the first-stage retrieval.
Table 4: Comparison the impact of using different MLLMs into the SceneRanker. Due to varying MLLM capabilities in processing structured prompts, we adopt Naive Prompt across all models to ensure fair comparison.
Model Parameters R@1 R@5 R@10 mAP
IRRA (ViT-L/14) 63.37 84.54 88.67 72.71
QwenVL2.5 bai2025qwen2 3 Billion 54.57 78.23 88.67 65.04
8 Billion 75.07 86.34 88.67 80.45
MiniCPM-V-2_6 yao2024minicpm 8 Billion 71.33 84.66 88.67 77.78
InternVL2.5 chen2024expanding 1 Billion 3.01 80.14 88.67 26.71
2 Billion 44.14 75.32 88.67 56.74
4 Billion 75.21 86.97 88.67 80.78
8 Billion 77.55 87.42 88.67 82.33
Different Prompt Design. We compare three strategies designed to enhance spatial grounding of pedestrian-specific region information in scene-aware multimodal prompts for person retrieval:
Prompt Design R@1 R@5 R@10 mAP 63.37 84.54 88.67 72.71 Naive Prompt 77.55 87.42 88.67 82.33 Box Overlay Prompt 76.69 87.29 88.67 81.85 Box Embedded Prompt 78.34 87.54 88.67 82.86
Table 5: Impact of visual prompt design on the performance of SceneRanker. IRRA(ViT-L/14) is adopted as the baseline for comparison, without prompt design.
1) Naive Prompt (NP), which provides only the textual description and the full-scene image without explicit region information; 2) Box Overlay Prompt (BOP), which visually highlights the candidate region with a red rectangle on the full-scene image; 3) Box Embedded Prompt (BEP), which includes the bounding box coordinates of the candidate region in the multimodal prompt using <box> tags. The complete designs of the three prompt strategies are provided in Appendix A.5. As shown in Table 5, incorporating box embedded prompt into the proposed SceneRanker outperforms naive prompt by 0.79% at R@1 and 0.53% at mAP, confirming that explicitly encoding region locations strengthens the model’s spatial grounding. Compared to the box overlay prompt, box embedded prompt achieves an even greater improvement of 1.65% at R@1 and 1.01% at mAP, suggesting that structured coordinate inputs are more effective than visual overlays. Notably, BOP even play worse than NP by 0.86% at R@1 and 0.48% at mAP, indicating that direct visual highlighting may introduce visual distractions that interfere with the model’s scene understanding.
Different Multimodal Large Language Models. We evaluate the effect of integrating different 8B-scale MLLMs into the proposed SceneRanker, using the same naive prompt for fair comparison. As shown in Table 4, all three models, Qwen-VL 2.5, MiniCPM-V, and InternVL 2.5, consistently improve over the IRRA baseline, with respective R@1 gains of 11.7%, +7.96%, and +14.18%. In addition, we observe a strong correlation between model capacity and performance. Taking InternVL 2.5 as an example, scaling from 4B to 8B brings further improvement, with R@1 increasing from 75.21% to 77.55%. In contrast, smaller models below 4B such as InternVL 1B and 2B show significantly lower performance, indicating insufficient multimodal understanding. These findings suggest that various MLLMs are well-suited for SceneRanker, and its performance benefits from stronger vision-language reasoning capabilities. Importantly, this holds true regardless of the specific architecture, provided the model has sufficient capacity.
Different Candidate Sizes. In the proposed two-stage framework SA-Person, the candidate set produced by the appearance-based retrieval stage serves as the input for the SceneRanker to perform scene-aware reranking. The size of this set ($K$) directly affects the trade-off between including the correct match and introducing noisy distractors. To better understand this trade-off and determine an optimal value for ($K$), we conduct a parameter analysis to systematically
Figure 3: Impact of candidate size.
investigate the impact of different candidate set sizes on the final retrieval performance. As shown in Figure  3, the performance of SceneRanker exhibits a trend of first increasing and then decreasing as the candidate size changes. The best performance is achieved when the candidate size is set to 10. Smaller candidate sets are more likely to exclude the ground-truth instance, while excessively large candidate sets introduce more irrelevant or noisy candidates. Both scenarios negatively impact performance. These results indicate that a moderate candidate pool is optimal, as it balances retaining the ground-truth instance while minimizing noise, ultimately enhancing retrieval accuracy.
4.5 Qualitative Results
To further validate the effectiveness of SA-Person, we present a qualitative comparison in Figure 4. IRRA focuses solely on pedestrian appearance, retrieving candidates with similar coarse attributes, but fails to utilize scene information and thus struggles to distinguish individuals in different contexts. LongCLIP, on the other hand, leverages scene-level information to retrieve images consistent with the described environment, but lacks fine-grained attention to pedestrian appearance and often returns individuals who differ significantly from the target identity. In comparison, SA-Person first retrieves candidates based on pedestrian appearance and then employs an MLLM to understand scene context and re-rank these candidates accordingly. This approach enables SA-Person to more accurately align with both the appearance and context described in the query, leading to significantly improved retrieval accuracy over existing baselines.
Figure 4: Qualitative comparison of retrieval results across three methods. The leftmost column shows the textual query and its corresponding ground-truth image. The top-6 retrieval results are presented from top to bottom for IRRA (ViT-L/14), LongCLIP (ViT-L/14), and SA-Person (ours).
5 Conclusion
In this paper, we tackle the challenging task of person retrieval by leveraging multimodal information from both pedestrian appearance and global scene context. Specifically, we make two key contributions. First, we present ScenePerson-13W, the first large-scale dataset to offer comprehensive multimodal annotations, integrating fine-grained pedestrian appearance descriptions with rich scene-level contextual information within full-scene images. Second, we propose SceneRanker, a novel scene-aware re-ranking method that exploits multimodal cues from the entire scene to effectively refine retrieval results. Building upon these contributions, we develop SA-Person, a two-stage framework that initially grounds discriminative appearance features in cropped pedestrian regions and subsequently applies SceneRanker, a training-free multimodal large language model—to perform scene-aware ranking based on holistic image context. Extensive experimental results demonstrate that SA-Person significantly outperforms existing baselines, underscoring the critical importance of integrating both local appearance cues and global scene context for robust and effective text-based person retrieval in complex real-world scenarios.
Limitations. While our work advances scene-aware person retrieval, it has two main limitations. First, the query descriptions in ScenePerson-13W are generated by MLLM, which may slightly deviate from natural human expression in retrieval scenarios. Second, our SceneRanker builds on off-the-shelf MLLMs without task-specific training, leaving room for improvement through fine-tuning to enhance alignment with retrieval intent and strengthen scene context understanding.
Appendix A Appendix of SA-Person: Text-Based Person Retrieval with Scene-aware Re-ranking
In the appendix, we provide the following materials:
  • Section A.1 provides a discussion of related work on text-based person retrieval and multimodal large language models.
  • Section A.2 analyzes the broader impact of scene-aware person retrieval, addressing societal benefits and ethical considerations.
  • Section A.3 details the construction and description generation processes for the SCENEPERSON-13W dataset.
  • Section A.4 presents supplemental results based on varisou retrieval methods and prompt designs, and visualizations of SceneRanker’s performance in Section 4.4 and 4.5 in the main paper.
  • Section A.5 describes prompt designs and instructional templates for SceneRanker in Section 4.4 in the main paper.
A.1 Related works
Text-Based Person Retrieval. Text-Based Person Retrieval (TBPR) aims to retrieve specific individuals based on natural language descriptions, typically from a gallery of cropped person images [45, 19, 55, 26, 25]. These methods focus on fine-grained cross-modal alignment between visual appearances and textual attributes, such as clothing, gender, or accessories [34, 55, 37, 38, 51]. With the advent of vision-language models like CLIP [31], TBPR performance has significantly improved. For instance, IRRA [15] introduces a multi-modal interaction encoder for implicit relational reasoning, and TBPS-CLIP [5] proposes a simple yet strong baseline through enhanced training strategies. To address noise in cross-modal data, RaSa [3] proposes sensitivity-aware and relation-aware representation learning for robustness against weakly matched pairs; MARS [10] incorporates visual reconstruction and attribute-sensitive losses to mitigate inter-identity noise and intra-identity variations; RDE [30] introduces dual embeddings to handle under-correlated or mislabeled image-text pairs under noisy conditions. Beyond cropped-image retrieval, some methods [50, 35, 42] retrieve individuals directly from unprocessed full-scene images, where individuals need to be located and identified in complex visual environments. Specifically, Zhang et al. [50] integrates textual cues into region proposals, enabling text descriptions to guide the detection process. Su et al. [35] introduces a memory-aided coarse-to-fine alignment strategy, using memory banks to enhance cross-modal contrastive learning. Yan et al. [42] explores multimodal fusion by combining textual and visual features with precise segmentation techniques to boost retrieval performance. While these methods often overlook the rich contextual information present in the surrounding scene, which could offer valuable clues for more accurate and robust person retrieval. Our SA-Person framework distinguishes itself by explicitly integrating scene context with appearance-based retrieval, leveraging textual descriptions of both individuals and their surroundings to enhance accuracy and disambiguate similar-looking candidates.
Multimodal Large Language Models for Scene and Region-Level Reasoning. Recent advances in Multimodal Large Language Models (MLLMs) enabled powerful visual reasoning across complex scenes. Current efforts primarily progresses along two directions: comprehensive scene understanding using global image-text inputs [7, 8, 2, 36, 28, 11], and fine-grained region-level grounding [13, 54, 29, 6, 27, 32]. For comprehensive scene understanding, recent works [28, 11] enable MLLMs to explicitly represent structured and compositional semantics. For instance, Mitra et al. [28] proposes generating scene graphs via MLLMs and incorporating them into prompts, facilitating compositional reasoning without relying on annotated scene graphs; MLLM-SUL [11] introduces a dual-branch encoder and a fine-tuned LLaMA [36] model to jointly perform scene description in traffic scenarios. To support region-level grounding, recent studies [13, 54, 6] enhance MLLMs with explicit spatial conditioning. For example, RegionGPT [13] and RegionBLIP [54] introduce region-wise feature pooling within bounding boxes or masks to extract localized representations for improved spatial reasoning; KOSMOS-2 [29] and Shikra [6] propose to encode spatial positions as discrete tokens or numeric coordinates to support referential understanding; GLaMM [32] facilitates grounded interaction through joint generation of text and segmentation masks. In this work, we explore the application of MLLMs to fine-grained person retrieval in complex scenes, where both textual descriptions and full-scene images are jointly leveraged to assess and rank visual relevance.
A.2 Broader impacts
The SA-Person framework and SCENEPERSON-13W dataset advance text-based person retrieval, improving public safety, smart city accessibility, and multimodal AI research. However, their surveillance applications risk privacy violations and biases from YouTube-derived data. To mitigate, restricting access with ethical guidelines ensures only responsible, research-focused use, preventing misuse by malicious actors and aligning with legal standards. Enhancing robustness through fine-tuning improves model accuracy across diverse scenarios, reducing errors and increasing reliability, thus maximizing societal benefits while ensuring ethical deployment.
Figure 5: Overview of person retrieval limitation and scene-aware insights. (a) Illustration of appearance-based retrieval limitation, where multiple candidates match the query “A girl wearing a green and white striped shirt is smiling” due to similar appearances. (b) Illustration of scene-aware reranking, ranking candidate images by aligning scene context like “located in a restaurant, near a window with red text” within the query. (c) Wordcloud of contextual elements in ScenePerson-13W, highlighting diverse scene descriptors. (d) Examples from the ScenePerson-13W dataset, showcasing detailed scene and relational descriptions for person retrieval.
Figure 6: Example instances from the ScenePerson-13W dataset. Descriptions in green text detail pedestrian appearance, while those in blue text describe their interactions within complex scenes. To clearly identify the target pedestrian associated with each textual description in scenes containing multiple pedestrians, the target individual is highlighted with a red bounding box in these visualizations.
A.3 More Details of ScenePerson-13W
The ScenePerson-13W dataset addresses critical challenges in scene-aware person retrieval by integrating detailed pedestrian appearance with rich scene context, overcoming limitations of appearance-only methods, as illustrated in Figure 5. This section provides additional details on the construction pipeline of the SCENEPERSON-13W dataset, with a focus on specific methodologies and confidence thresholds to ensure reproducibility and clarity.
Pedestrian Detection and Tracking. Using the YOLOv11x, we detect and track pedestrians in the QVHighlights video dataset [17] with a confidence threshold of 0.6 to balance precision and recall.
Completeness Filtering. To ensure quality, incomplete detections are filtered using MediaPipe’s pose estimation [1] with a minimum detection confidence of 0.65. Head detection requires at least one keypoint (e.g., nose, eyes) with visibility over 0.4, and shoulders need visibility over 0.6 for at least one shoulder. Only images with a head and one shoulder are retained.
Deduplication. Redundancy is reduced using CLIP [31] to compute cosine similarity, normalized to the range [0, 1], for cropped pedestrian regions and full-scene images. Images with similarity score exceeding 0.91 are deemed redundant, and one instance is kept, yielding 131,438 unique pedestrian instances for diversity.
Testing Set Refinement. To create a fair and robust testing set of 10,000 images, we address redundancy from frame-based tracking, where pedestrians appear across consecutive frames with minor variations. We apply CLIP-based visual clustering using the clip-vit-base-patch32 [31] model, with a similarity threshold of 0.88. Images are clustered via agglomerative clustering with average linkage, and one representative image per cluster is selected. This deduplicated testing set balances diversity and manageability, preserving complex scene and identity distributions for meaningful retrieval evaluation.
Description Generation. Textual descriptions for each of the 131,438 pedestrian instances are generated using InternVL2.5-8B [8]. The model is prompted with a full-scene image, highlighting the target pedestrian with a red bounding box, using the following prompt:
Generate a description of the individual within the red bounding box and the connection with the surroundings, in a style similar to the template: template. Ensure that the description is concise within 50 words and accurate. Do not include text details or labels, and strictly avoid any punctuation other than commas and periods.
The template guides the model to describe: (1) the individual’s appearance, (2) their location in the scene, and (3) relationships with surroundings. Specifically, the template is defined as:
A [man/woman/girl/boy] wearing [describe basic clothing, notable accessories, and physical attributes] is [describe their action or posture]. This person is located [describe their location, highlighting any landmarks, structures, or recognizable surroundings]. Next to/to the left of/to the right of/behind/in front of/under/above/around [the man/woman/etc.], there is/are [briefly describe nearby people, objects, or structures, such as another person in red holding something, a group of people talking, or a pedestrian bridge].
This structured approach ensures descriptions capture both pedestrian details and contextual cues, as illustrated in Figure 6.
Dataset Examples. More examples of ScenePerson-13W are shown in Figure 6. showcasing the dataset’s rich scene and relational descriptions. These examples highlight the dataset’s ability to provide fine-grained appearance details and diverse scene contexts, addressing challenges like those in Figure 5, where scene-aware insights disambiguate similar-looking individuals.
A.4 Experiments and Visualisation
In this section, we provide deeper insights into our experimental results, ablation findings, and qualitative visualizations.
A.4.1 SceneRanker Performance Across Retrieval Methods and Prompt Designs
Table 6: Performance of SceneRanker with various text-based person retrieval methods under different prompt designs. Red numbers indicate the improvement brought by each prompt design.
     Methods      R@1      R@5      R@10      mAP
     TBPS-CLIP [5]      50.29      75.24      81.58      61.54
      +NP      $71.41_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+21.12}}$      $80.19_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+4.95}}$      $81.58_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+0.00}}$      $76.03_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+14.49}}$
      +BOP      $70.77_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+20.48}}$      $80.16_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+4.92}}$      $81.58_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+0.00}}$      $75.69_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+14.15}}$
      +BEP      $\textbf{72.25}_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+21.96}}$      $\textbf{80.23}_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+4.99}}$      $\textbf{81.58}_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+0.00}}$      $\textbf{76.54}_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+15.00}}$
     IRRA(ViT-B/16) [15]      60.57      83.35      87.86      70.65
      +NP      $76.80_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+16.23}}$      $86.41_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+3.06}}$      $87.86_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+0.00}}$      $81.53_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+10.88}}$
      +BOP      $75.78_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+15.21}}$      $86.39_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+3.04}}$      $87.86_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+0.00}}$      $80.99_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+10.34}}$
      +BEP      $\textbf{77.42}_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+16.85}}$      $\textbf{86.61}_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+3.26}}$      $\textbf{87.86}_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+0.00}}$      $\textbf{81.96}_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+11.31}}$
     IRRA(ViT-L/14) [15]      63.37      84.54      88.67      72.71
      +NP      ${77.55}_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+14.18}}$      ${87.42}_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+2.88}}$      ${88.67}_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+0.00}}$      ${82.33}_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+9.62}}$
      +BOP      ${76.69}_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+13.32}}$      ${87.29}_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+2.75}}$      ${88.67}_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+0.00}}$      ${81.85}_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+9.14}}$
      +BEP      $\textbf{78.34}_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+14.97}}$      $\textbf{87.54}_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+3.00}}$      $\textbf{88.67}_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+0.00}}$      $\textbf{82.86}_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+10.15}}$
     RDE [30]      61.26      83.50      88.23      71.17
      +NP      $77.13_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+15.87}}$      $86.83_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+3.33}}$      $88.23_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+0.00}}$      $81.87_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+10.70}}$
      +BOP      $76.29_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+15.03}}$      $86.88_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+3.38}}$      $88.23_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+0.00}}$      $81.46_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+10.29}}$
      +BEP      $\textbf{77.58}_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+16.32}}$      $\textbf{86.84}_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+3.34}}$      $\textbf{88.23}_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+0.00}}$      $\textbf{82.18}_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+11.01}}$
     RaSa [3]      62.51      83.83      88.09      72.05
      +NP      $77.04_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+14.53}}$      $86.66_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+2.83}}$      $88.09_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+0.00}}$      $81.76_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+9.71}}$
      +BOP      $76.24_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+13.73}}$      $86.74_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+2.91}}$      $88.09_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+0.00}}$      $81.29_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+9.24}}$
      +BEP      $\textbf{77.84}_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+15.33}}$      $\textbf{86.83}_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+3.00}}$      $\textbf{88.09}_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+0.00}}$      $\textbf{82.27}_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+10.22}}$
     MARS [10]      62.93      84.26      88.65      72.43
      +NP      $77.26_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+14.33}}$      $87.30_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+3.04}}$      $88.65_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+0.00}}$      $82.10_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+9.67}}$
      +BOP      $76.57_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+13.64}}$      $87.14_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+2.88}}$      $88.65_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+0.00}}$      $81.69_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+9.26}}$
      +BEP      $\textbf{78.32}_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+15.39}}$      $\textbf{87.39}_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+3.13}}$      $\textbf{88.65}_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+0.00}}$      $\textbf{82.77}_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+10.34}}$
To further assess SceneRanker’s generalizability, we evaluate its performance when integrated with various text-based person retrieval methods, each paired with three prompt designs: Naive Prompt (NP), Box Overlay Prompt (BOP), and Box Embedded Prompt (BEP). Results are presented in Table 6. SceneRanker significantly boosts performance for all methods, with R@1 improvements exceeding above 14.97% when using BEP, underscoring its compatibility and effectiveness in refining appearance-based retrieval. Across all methods, BEP consistently outperforms NP and BOP, achieving the highest R@1, suggesting that embedding bounding box coordinates enhances spatial grounding, enabling better alignment with textual descriptions.
A.4.2 Visualization of Retrieval Results
To illustrate the impact of SceneRanker, we present a qualitative comparison of retrieval results before and after applying SceneRanker, as shown in Figure 7.
Figure 7: Qualitative Comparison of Retrieval Results with and without SceneRanker. The leftmost column shows the query and ground-truth image. The green frame contains the retrieval results from the appearance-based stage before SceneRanker re-ranking, while the blue frame contains the retrieval results after SceneRanker re-ranking.
A.5 SceneRanker Prompt Design and Templates
This subsection details the instructional prompt templates used in the SceneRanker ablation studies. SceneRanker performs reranking without task-specific fine-tuning, using a multimodal prompt comprising a visual prompt $\mathbf{P}_{v}$ and an instructional textual prompt. Within each prompt strategy, $\mathbf{P}_{v}$ is fixed, but the instructional prompt varies to enhance spatial grounding, with multiple prompts tested. We describe: (1) the instructional prompt templates for three strategies evaluated for enhancing spatial grounding in SceneRanker, each in a dedicated table, and (2) the model-specific instructional prompt templates designed for Qwen-VL 2.5 [2], MiniCPM-V-2_6 [46], and InternVL 2.5 [8] in the MLLMs ablation study, tailored to each model’s capabilities while ensuring fair comparison. All prompts are designed for use with the InternVL2.5-8B model [8] unless otherwise specified.
A.5.1 Scene-aware Multimodal Prompt Templates
For each prompt strategy, the visual prompt $\mathbf{P}_{v}$ is fixed, but multiple instructional prompts were tested to optimize spatial grounding of pedestrian-specific region information. Below, we specify $\mathbf{P}_{v}$ for each strategy, describe the instructional prompt variations for reproducibility, and present examples in dedicated tables, with the final adopted prompt in bold.
[…]
"For the text description: ’text’, order the images based on how accurately they reflect overall context in the text, starting with the most faithful match. Please output only a Python list format like this: [4, 2, 8, 1, 5, 3, 6, 9, 10, 7]." "For the description: ‘$T$’, 10 images are provided. Each image includes one person who is the intended candidate subject described in the text. Order the images based on how accurately they reflect overall context in the text, starting with the most faithful match. Please output only a Python list format like this: [4, 2, 8, 1, 5, 3, 6, 9, 10, 7]." "For the description: ‘$T$’, 10 images are provided. In each image, one person is implicitly assumed to be the potential match for the individual described. Order the images based on how accurately they reflect overall context in the text, starting with the most faithful match. Please output only a Python list format like this: [4, 2, 8, 1, 5, 3, 6, 9, 10, 7]." "For the given description: ‘$T$’, analyze the 10 provided candidate images. Each image contains one person implicitly corresponding to a potential subject referenced in the description. Evaluate how accurately each candidate aligns with the contextual details, visual cues, and semantic elements in the text description. Rank these images in descending order of relevance and contextual fidelity, with the most accurate match first. Return only a Python list of indices in the format: [4, 2, 8, 1, 5, 3, 6, 9, 10, 7], where each number represents the original position of the image in the candidate set." "Given the description: ‘$T$’, 10 images are provided, each containing a person implicitly regarded as the candidate subject described in the text. Reorder the images based on how well they match the description, considering both the assumed subject’s appearance and the surrounding context. Provide the output as a Python list in descending order of relevance, formatted like this: [4, 2, 8, 1, 5, 3, 6, 9, 10, 7]."
Table 7: Naive Prompt (NP) instructional prompt templates for SceneRanker.
"For the text description: ‘$T$’, order the images based on how accurately they reflect overall context in the text, starting with the most faithful match. Please output only a Python list format like this: [4, 2, 8, 1, 5, 3, 6, 9, 10, 7]." "For the description: ‘$T$’, 10 images are provided. Each image includes one person highlighted with a bounding box, representing the candidate subject mentioned in the text. Order the images based on how accurately they reflect overall context in the text, starting with the most faithful match. Please output only a Python list format like this: [4, 2, 8, 1, 5, 3, 6, 9, 10, 7]." "For the description: ‘$T$’, 10 images are provided. In each image, a person is marked with a bounding box as the potential match for the individual described. Order the images based on how accurately they reflect overall context in the text, starting with the most faithful match. Please output only a Python list format like this: [4, 2, 8, 1, 5, 3, 6, 9, 10, 7]." "For the given description: ‘$T$’, analyze the 10 provided candidate images. Each image contains one person highlighted with a bounding box, representing a potential subject referenced in the description. Evaluate how accurately each candidate aligns with the contextual details, visual cues, and semantic elements in the text description. Rank these images in descending order of relevance and contextual fidelity, with the most accurate match first. Return only a Python list of indices in the format: [4, 2, 8, 1, 5, 3, 6, 9, 10, 7], where each number represents the original position of the image in the candidate set." "Given the description: ‘$T$’, 10 images are provided, each containing a person highlighted with a bounding box, representing the candidate subject described in the text. Reorder the images based on how well they match the description, considering both the highlighted subject’s appearance and the surrounding context. Provide the output as a Python list in descending order of relevance, formatted like this: [4, 2, 8, 1, 5, 3, 6, 9, 10, 7]."
Table 8: Box Overlay Prompt (BOP) instructional prompt templates for SceneRanker.
"For the text description: ‘$T$’, order the images based on how accurately they reflect overall context in the text, starting with the most faithful match. Please output only a Python list format like this: [4, 2, 8, 1, 5, 3, 6, 9, 10, 7]." "For the text description: ‘$T$’, 10 images are provided. Each image includes one person marked by a bounding box (given as coordinates in <box>). This bounding box indicates the specific person selected as a potential match for the description. Please rank the images based on how accurately the boxed person and their surrounding scene match the description, starting with the most accurate match. Please output only a Python list format like this: [4, 2, 8, 1, 5, 3, 6, 9, 10, 7]." "For the text description: ‘$T$’, 10 images are provided. Each image includes a bounding box, which indicates the specific person selected as a potential match for the description, and its coordinates are shown as <box> for each image. Please rank the images based on how accurately the boxed person and their surrounding scene match the description, starting with the most accurate match. Please output only a Python list format like this: [4, 2, 8, 1, 5, 3, 6, 9, 10, 7]." "For the description: ‘$T$’, 10 images are provided. In each image, one person is implicitly assumed to be the potential match for the individual described. Order the images based on how accurately they reflect overall context in the text, starting with the most faithful match. Please output only a Python list format like this: [4, 2, 8, 1, 5, 3, 6, 9, 10, 7]." "For the description: ‘$T$’, 10 images are provided. Each image includes one person who is the intended candidate subject described in the text. Order the images based on how accurately they reflect overall context in the text, starting with the most faithful match. Please output only a Python list format like this: [4, 2, 8, 1, 5, 3, 6, 9, 10, 7]."
Table 9: Box Embedded Prompt (BEP) instructional prompt templates for SceneRanker.
[…]
A.5.2 Model-Specific Instructional Prompt Templates
For the MLLM ablation study, SceneRanker is evaluated with Qwen-VL 2.5, MiniCPM-V-2_6, and InternVL 2.5, each using a tailored instructional prompt template paired with plain full-scene images. These prompts are designed to leverage each model’s strengths while ensuring fair comparison. Their respective instructional prompt templates are provided in Table 10.
Qwen-VL 2.5: "For the text description: ‘$T$’, analyze the 10 candidate images and rank them from most to least consistent with the text. Output only a Python list containing all integers from 1 to 10, with no duplicates or omissions, like [4, 2, 8, 1, 5, 3, 6, 9, 10, 7]." MiniCPM-V: "Considering the full context of the description ‘$T$’, rank the provided images from best to worst match. Output only a Python list containing numbers from 1 to 10." InternVL 2.5: "For the text description: ‘$T$’, order the images based on how accurately they reflect overall context in the text, starting with the most faithful match. Please output only a Python list format like this: [4, 2, 8, 1, 5, 3, 6, 9, 10, 7]."
Table 10: Instructional Prompt Templates for Comparative Evaluation of Multimodal Large Language Models.

### v3

@@
 Abstract
-Text-based person retrieval aims to identify a target individual from a gallery of images based on a natural language description.
-It presents a significant challenge due to the complexity of real-world scenes and the ambiguity of appearance-related descriptions.
-Existing methods primarily emphasize appearance-based cross-modal retrieval, often neglecting the contextual information embedded within the scene, which can offer valuable complementary insights for retrieval.
-To address this, we introduce ScenePerson-13W, a large-scale dataset featuring over 100,000 scenes with rich annotations covering both pedestrian appearance and environmental cues.
-Based on this, we propose SA-Person, a two-stage retrieval framework.
-In the first stage, it performs discriminative appearance grounding by aligning textual cues with pedestrian-specific regions.
-In the second stage, it introduces SceneRanker, a training-free, scene-aware re-ranking method leveraging multimodal large language models to jointly reason over pedestrian appearance and the global scene context.
-Experiments on ScenePerson-13W validate the effectiveness of our framework in challenging scene-level retrieval scenarios.
-The code and dataset will be made publicly available.
-00footnotetext: * These authors contributed equally to this work.00footnotetext: †Corresponding author.
-1 Introduction
-Person retrieval galiyawala2021person aims to identify a target individual within the image galleries based on a query, playing a vital role in applications like video surveillance and human-centric artificial intelligent systems li2018richly ; galiyawala2021person ; li2023dcel .
-Early efforts focus on image-based paradigms zhang2016learning ; liu2018pose ; sarfraz2018pose ; xiao2017joint ; he2018end ; yan2021anchor ; cao2022pstr ; yan2022exploring using cropped images as queries, and then shift toward text-based retrieval li2017person ; yu2019cross ; liu2023survey due to the accessibility of textual descriptions.
-Advances in vision-language pretraining (VLP) models, such as CLIP radford2021learning , also accelerate its development yang2023towards ; li2024adaptive ; zuo2024ufinebench ; luo2025graph ; liu2025dm ; jiang2023cross ; bai2023rasa ; cao2024empirical ; ergasti2024mars ; qin2024noisy .
-However, text-based methods rely solely on cropped images and textual descriptions of appearance, limiting their effectiveness in real-world scenarios.
-Recent methods zhang2023text ; su2024maca ; yan2025fusionsegreid extend retrieval to full-scene images but still focus narrowly on appearance, neglecting critical environmental scene cues.
-This limitation poses challenges in complex scenarios with similar-looking individuals, or the same person under varying contexts.
-Incorporating contextual scene semantics alongside individual appearance is essential for accurate and robust person retrieval.
-This work focuses on scene-aware text-based person retrieval.
+Text-based person retrieval aims to identify a target individual from an image gallery using a natural language description.
+Existing methods primarily focus on appearance-driven cross-modal retrieval, yet face significant challenges due to the visual complexity of scenes and the inherent ambiguity of textual descriptions.
+The contextual information, such as landmarks and relational cues, provides complementary cues that can offer valuable complementary insights for retrieval, but remains underexploited in current approaches.
+Motivated by this limitation, we propose a novel paradigm: scene-aware text-based person retrieval, which explicitly integrates both individual appearance and global scene context to improve retrieval accuracy.
+To support this, we first introduce ScenePerson-13W, a large-scale benchmark dataset comprising over 100,000 real-world scenes with rich annotations encompassing both pedestrian attributes and scene context.
+Based on this dataset, we further present SA-Person, a two-stage retrieval framework.
+In the first stage, SA-Person performs discriminative appearance grounding by aligning textual descriptions with pedestrian-specific regions.
+In the second stage, it introduces SceneRanker, a training-free, scene-aware re-ranking module that refines retrieval results by jointly reasoning over pedestrian appearance and the global scene context.
+Extensive experiments on ScenePerson-13W and existing benchmarks demonstrate the effectiveness of our proposed SA-Person.
+Both the dataset and code will be publicly released to facilitate future research.
+I Introduction
+Person retrieval [galiyawala2021person] aims to identify a target individual within the image galleries based on a query, playing a vital role in applications like video surveillance and human-centric artificial intelligent systems [li2018richly, galiyawala2021person, li2023dcel].
+Early efforts focus on image-based paradigms [zhang2016learning, liu2018pose, sarfraz2018pose, xiao2017joint, he2018end, yan2021anchor, cao2022pstr, yan2022exploring] using cropped images as queries, and then shift toward text-based retrieval [li2017person, yu2019cross, liu2023survey] due to the accessibility of textual descriptions.
+Advances in large-scale Vision-Language Pretraining (VLP) models, such as CLIP [radford2021learning], also accelerate the development of text-based person retrieval [yang2023towards, li2024adaptive, zuo2024ufinebench, luo2025graph, liu2025dm, jiang2023cross, bai2023rasa, cao2024empirical, ergasti2024mars, qin2024noisy].
+However, methods for text-based person retrieval rely solely on cropped images and appearance-focused textual descriptions, limiting their effectiveness in real-world scenarios.
+Recent methods [zhang2023text, su2024maca, yan2025fusionsegreid] extend retrieval to full-scene images but still focus narrowly on appearance, neglecting critical scene cues.
+This limitation poses challenges in complex scenarios with similar-looking individuals, or the same person under varying contexts, as illustrated in Figure 1(a).
+In such cases, appearance-only retrieval often produces ambiguous or incorrect matches.
+By contrast, human descriptions about a target person are rarely restricted to appearance; instead, they naturally incorporate auxiliary contextual cues that complement visual traits, such as surrounding environments, landmark buildings, and co-occurring entities.
+As illustrated in Figure 1(b), incorporating scene context substantially enhance disambiguation beyond what appearance features alone can offer.
+To quantify the potential of scene context, we perform a preliminary analysis on the CUHK-SYSU dataset augmented with MLLM-generated descriptions.
+As shown in Figure 1(c), MLLMs with context-enriched text can achieve retrieval performance beyond the limits of fine-tuned domain-specific models.
+A comprehensive discussion is provided in Section III.
+Motivated by this observation, we focus on scene-aware text-based person retrieval, which explicitly incorporates both individual appearance and global scene context for more accurate and robust retrieval.
 The first challenge is the lack of a dedicated dataset for this task.
@@
 It includes over $131,438$ images from $12,562$ YouTube video sequences, featuring diverse individuals across varying crowd densities, camera angles, and environments.
-Unlike existing datasets that focus on cropped images or weakly grounded captions, ScenePerson-13W provides rich image-text pairs with detailed descriptions encompassing both visual appearance and contextual scene information, which promotes text-to-person alignment.
+As summarized in Table II, existing datasets[zheng2015scalable, li2014deepreid, xiao2017joint, zheng2017person, li2017person, ding2021semantically, zhang2023text] typically focus on cropped images or weakly grounded captions, ScenePerson-13W provides rich image-text pairs with detailed descriptions encompassing both visual appearance and scene context, which promotes text-to-person alignment.
 Building on ScenePerson-13W, we propose SA-Person, a two-stage framework for scene-aware person retrieval.
 In the first stage, appearance-related information is extracted from textual descriptions to identify candidate pedestrians via cross-modal matching, reducing the retrieval space with discriminative appearance features.
-In the second stage, we introduce SceneRanker, a training-free, scene-aware re-ranking module powered by MLLM.
-By employing structured multimodal prompts, SceneRanker enables holistic reasoning over appearance and global scene context, allowing effective disambiguation of visually similar individuals and significantly improving retrieval performance.
+In the second stage, we introduce SceneRanker, a training-free, scene-aware re-ranking module that refines retrieval results by jointly reasoning over pedestrian appearance and global scene context, enabling holistic understanding to effectively disambiguate visually similar individuals and significantly improve retrieval performance.
+Figure 1: Overview of person retrieval challenges and scene-aware insights.
+(a) Illustration of appearance-based retrieval limitations, where multiple candidates match the text A girl wearing a green and white striped shirt is smiling due to similar appearances.
+(b) Demonstration of scene-aware reranking, ranking candidate images by aligning scene context like located in a restaurant, near a window with red text with the text.
+(c) Retrieval performance across three progressive input configurations on the augmented CUHK-SYSU dataset.
+The three configurations are: (1) Cropped + App (cropped image with appearance-only text), (2) Full + App (full image with appearance-only text), and (3) Full + Context (full image with context-enriched text).
 In summary, our main contributions are as follows:
-  • We introduce ScenePerson-13W, a new large-scale dataset for scene-aware text-based person retrieval, with compositional descriptions that include both pedestrian appearance and global scene context, alleviating the limitations of prior datasets.
+  1. 1)
+    We introduce ScenePerson-13W, the first dataset enabling a new retrieval paradigm of scene-aware text-based person retrieval by providing compositional descriptions that include both pedestrian appearance and global scene context, alleviating the limitations of prior datasets.
+  2. 2)
     We present SA-Person, a two-stage retrieval framework that combines discriminative appearance grounding with holistic scene understanding, enabling robust retrieval from large-scale full-scene image databases.
-  • We propose SceneRanker, the first re-ranking strategy for text-based person retrieval powered by MLLM, which follows a training-free pattern and incorporates visual grounding and contextual reasoning to improve retrieval accuracy in complex scenes.
-  • Extensive experiments on ScenePerson-13W demonstrate the effectiveness and generalizability of our proposed SA-Person across various retrieval settings and model backbones.
+  3. 3)
+    We propose SceneRanker, the first re-ranking strategy for text-based person retrieval that can be seamlessly integrated as a plug-and-play module, incorporating visual grounding and contextual reasoning to disambiguate and enhanace retrieval performance in complex scenes.
+  4. 4)
+    Extensive experiments on ScenePerson-13W and diverse benchmarks demonstrate the effectiveness and generalizability of our proposed SA-Person across various retrieval settings and model backbones.
+Figure 2: Qualitative visualization of retrieval results across three progressive input configurations on the augmented CUHK-SYSU dataset.
+The blue frame contains the top candidates from the domain-specific model IRRA (ViT-L/14), and the orange frame contains the top candidates from the MLLM (InternVL-8B).
+The green-bordered image is the ground truth.
+II Related works
+II-A Text-Based Person Retrieval
+Text-Based Person Retrieval (TBPR) aims to retrieve specific individuals based on natural language descriptions, typically from a gallery of cropped person images [yang2023towards, li2024adaptive, zuo2024ufinebench, luo2025graph, liu2025dm].
+These methods focus on fine-grained cross-modal alignment between visual appearances and textual attributes, such as clothing, gender, or accessories [shu2022see, zuo2024ufinebench, wang2024fine, wei2024fine, zhao2025cross].
+With the advent of vision-language models like CLIP [radford2021learning], TBPR performance has significantly improved.
+For instance, IRRA [jiang2023cross] introduces a multi-modal interaction encoder for implicit relational reasoning, and TBPS-CLIP [cao2024empirical] proposed a simple yet strong baseline through enhanced training strategies.
+To address noise in cross-modal data, RaSa [bai2023rasa] proposed sensitivity-aware and relation-aware representation learning for robustness against weakly matched pairs; MARS [ergasti2024mars] incorporated visual reconstruction and attribute-sensitive losses to mitigate inter-identity noise and intra-identity variations; RDE [qin2024noisy] introduced dual embeddings to handle under-correlated or mislabeled image-text pairs under noisy conditions.
+Beyond cropped-image retrieval, some methods [zhang2023text, su2024maca, yan2025fusionsegreid] retrieval individuals directly from unprocessed full-scene images, where individuals must be located and identified in complex visual environments Specifically, Zhang et al.
+[zhang2023text] integrated textual cues into region proposals, enabling text descriptions to guide the detection process.
+Su et al.
+[su2024maca] introduced a memory-aided coarse-to-fine alignment strategy, using memory banks to enhance cross-modal contrastive learning.
+Yan et al.
+[yan2025fusionsegreid] explored multimodal fusion by combining textual and visual features with precise segmentation techniques to boost retrieval performance.
+While these methods often overlook the rich contextual information present in the surrounding scene, which could offer valuable clues for more accurate and robust person retrieval.
+II-B Multimodal Large Language Models for Scene and Region-Level Reasoning
+Recent advances in Multimodal Large Language Models (MLLMs) enabled powerful visual reasoning across complex scenes.
+Current efforts primarily progresses along two directions: comprehensive scene understanding using global image-text inputs [chen2024sharegpt4v, chen2024expanding, bai2025qwen2, touvron2023llama, mitra2024compositional, fan2024mllm], and fine-grained region-level grounding [guo2024regiongpt, zhou2023regionblip, peng2023kosmos, chen2023shikra, ma2024groma, rasheed2024glamm].
+For comprehensive scene understanding, recent works [mitra2024compositional, fan2024mllm] enables MLLMs to explicitly represent structured and compositional semantics.
+For instance, Mitra et al.
+[mitra2024compositional] proposed generating scene graphs via MLLMs and incorporating them into prompts, facilitating compositional reasoning without relying on annotated scene graphs; MLLM-SUL  [fan2024mllm] introduced a dual-branch encoder and a fine-tuned LLaMA [touvron2023llama] model to jointly perform scene description in traffic scenarios.
+To support region-level grounding, recent studies  [guo2024regiongpt, zhou2023regionblip, chen2023shikra] enhance MLLMs with explicit spatial conditioning.
+For example, RegionGPT [guo2024regiongpt] and RegionBLIP [zhou2023regionblip] have been proposed adopt region-wise feature pooling within bounding boxes or masks to extract localized representations for improved spatial reasoning; KOSMOS-2 [peng2023kosmos] and Shikra [chen2023shikra] encode spatial positions as discrete tokens or numeric coordinates to support referential understanding; GLaMM [rasheed2024glamm] enables grounded interaction through joint generation of text and segmentation masks.
+In this work, we combine the complementary strengths of domain-specific TBPR models and MLLMs for scene-aware text-based person retrieval, fusing efficient appearance alignment with holistic context reasoning to robustly disambiguate in complex scenes.
+TABLE I: Retrieval performance across three progressive input configurations on CUHK-SYSU dataset, augmented using MLLM-generated descriptions.
+Cropped Image Full Image Full Image
+& Appearance-only Text & Appearance-only Text & Context-enriched Text
+Model R@1 R@5 R@10 mAP R@1 R@5 R@10 mAP R@1 R@5 R@10 mAP
+Domain-specific TBPR models (zero-shot)
+RDE [qin2024noisy] (ViT-B/16) 40.62 79.19 84.18 57.34 21.17 50.09 62.28 35.19 31.82 69.58 81.32 48.53
+IRRA [jiang2023cross] (ViT-B/16) 41.62 79.45 84.73 58.27 23.23 53.47 66.09 37.76 35.78 74.68 85.39 53.04
+Domain-specific TBPR models (fine-tuned)
+IRRA [jiang2023cross] (ViT-B/16) 47.60 84.00 89.14 63.63 33.76 73.69 84.40 51.37 66.61 98.53 99.60 80.72
+RDE [qin2024noisy] (ViT-B/16) 48.73 83.78 88.88 63.93 35.12 73.47 82.97 51.98 67.23 98.31 99.60 81.05
+IRRA [jiang2023cross] (ViT-L/14) 50.57 85.54 89.36 66.01 36.81 76.66 86.13 54.63 71.41 98.86 99.93 83.56
+Multimodal large model (zero-shot)
+InternVL-8B [chen2024expanding] 23.38 40.92 50.79 31.03 46.32 80.37 92.13 59.39 85.27 97.53 98.90 90.63
+III Preliminary Analysis of Current Models on Text-Based Person Retrieval
+Although text-based person retrieval has made significant strides driven by vision-language pretraining models, existing methods [jiang2023cross, bai2023rasa, cao2024empirical, ergasti2024mars, qin2024noisy] predominantly rely on appearance descriptions and cropped pedestrian images, overlooking the potential value of scene context.
+To empirically substantiate this oversight and demonstrate the complementary strengths of domain-specific TBPR models and MLLMs in handling increasing scene complexity, which thereby motivating our two-stage SA-Person framework, we conduct a preliminary analysis on the augmented CUHK-SYSU [xiao2017joint].
+III-A Analysis Configuration
+Originally featuring full-scene images with annotated target pedestrian bounding boxes, we augment CUHK-SYSU with MLLM-generated textual descriptions, with two types per target: (1) appearance-only descriptions, capturing visual attributes such as clothing, accessories, and pose; and (2) context-enriched descriptions, augmenting these with contextual details that include surroundings, spatial relations and interactions.
+This enables a gradual incorporation of scene context, tracing the shift from appearance-focused to context-aided retrieval through three progressive input configurations:
+  1. 1)
+    Appearance-only text & cropped image: Relies solely on appearance-focused descriptions and the target’s cropped bounding box region, isolating appearance alignment.
+  2. 2)
+    Appearance-only text & full image: Pairs the same descriptions with complete scene images, introducing scene information without explicit contextual guidance.
+  3. 3)
+    Context-enriched text & full image: Combines augmented descriptions with full scenes, evaluating joint reasoning over appearance and environmental semantics.
+Evaluated models span two architectural paradigms: (1) state-of-the-art domain-specific TBPR models, including IRRA [jiang2023cross] and RDE [qin2024noisy], which use dual-encoder frameworks.
+For this analysis, we finetune these models on the augmented CUHK-SYSU dataset, adapting to each configuration’s input modality (e.g., holistic visuals for the last setting).
+We also evaluate their zero-shot performance using pre-trained weights from CUHK-PEDES.
+(2) Multimodal large language models (MLLMs), such as InternVL-8B [chen2024expanding], which support zero-shot retrieval without modality-specific adaptation.
+However, current MLLMs struggle with gallery-wide retrieval across vast image galleries, constrained by their bounded context capacities that restrict simultaneous handling of extensive images [niu2025chatreid].
+To address this, we transform the complex retrieval task into a multi-round scoring and ranking process through engineering enhancements.
+III-B Performance Analysis
+Table I presents results across the three progressive configurations, revealing critical insights into the role of scene context in TBPR and highlighting the complementary strengths of domain-specific TBPR models and MLLMs when handling scene complexity.
+In the appearance-only text & cropped image setting, domain-specific models dominate.
+Zero-shot IRRA pre-trained on CUHK-PEDES achieves 41.62% R@1, rising to 50.57% after fine-tuning, both surpassing zero-shot InternVL-8B’s 23.38% R@1.
+Their specialized architecture and extensive appearance-focused pretraining yield highly discriminative embeddings for clean, isolated regions.
+In contrast, zero-shot MLLMs lack domain alignment, producing coarser cross-modal representations.
+However, switching to appearance-only text & full image setting, full scenes without contextual support, domain-specific models exposes a sharp vulnerability.
+IRRA with ViT-L/14 drops from 50.57% to 36.81% R@1, and zero-shot domain models fall to $\sim$23%.
+Distracting environmental elements like cluttered backgrounds and occlusions overwhelm appearance-based embeddings lacking semantic safeguards.
+In contrast, zero-shot InternVL-8B rises to 46.32% R@1, leveraging inherent scene understanding to mitigate interference.
+Shifting to the context-enriched text & full image setting further underscores the value of scene context, where contextual augmentation markedly enhances retrieval accuracy across models.
+As shown in Figure 2, the contextual support enables refined disambiguation.
+Zero-shot IRRA yields only 35.78% R@1 due to limited generalization to context-enriched descriptions.
+Fine-tuning boosts R@1 to 71.41%, a 99% relative gain, yet still falls short of the zero-shot MLLM baseline, where InternVL-8B attains 85.27% R@1, demonstrating clear superior inherent generalization and seamless integration of local appearance with global elements like landmarks and relational cues.
+Domain-specific models, despite their improvements, cannot fully leverage these aspects.
+III-C Time Complexity Analysis
+For computational cost, MLLMs incur substantially higher expenses than domain-specific models due to the multi-round gallery retrieval process.
+Let $\mu_{s}$ and $\mu_{m}$ denote the per-image time units for domain-specific models and MLLMs, respectively, with $\mu_{m}\approx 2$ orders of magnitude larger than $\mu_{s}$ due to intensive multimodal processing.
+Domain-specific models require $O(N\mu_{s})$ for a gallery of size $N$.
+MLLMs demand an initial $O(N)$ traversal for match scores at $O(N\mu_{m})$ time, plus $O(M/B)$ batch ranking operations with per-batch time $T_{b}=\mu_{m}\times B$, where $M$ is the candidate set size and $B$ is the batch size.
+The total time thus simplifies to $O(N\mu_{m}+K\mu_{m})$.
+This yields an overall per-query time orders of magnitude higher than that of domain-specific models, scaling linearly with $N$ and limiting scalability for large galleries.
+In summary, this analysis reveals a key contrast: domain-specific models excel in straightforward appearance tasks but demonstrate poor zero-shot generalization and struggle with scene variations, while MLLMs deliver superior and robust holistic reasoning in context-rich environments at a higher computational cost.
+The widening performance gaps as scene context increases underscore the value of approaches that connect detailed region matching and broader reasoning, directly motivating our development of SCENEPERSON-13W and the two-stage SA-Person framework to harness these complementary strengths for robust, real-world TBPR.
 Figure 3: Overview of the ScenePerson-13W construction pipeline.
-The construction pipeline involves scene segmentation, pedestrian detection and tracking, completeness filtering, image deduplication, and textual description generation.
+The construction pipeline involves scene segmentation, pedestrian detection and tracking, completeness filtering, image deduplication, and description generation.
 For each retained pedestrian, a description is generated based on the full image with a highlighted target, capturing their appearance, spatial location, and relationships with surrounding elements.
-Table 1: Statistics of person retrieval datasets.
+TABLE II: Statistics of person retrieval datasets.
 SCENEPERSON-13W shows a substantial advancement over existing person retrieval datasets by combining fine-grained appearance details with rich global scene contexts.
   Dataset   Query Modality   Gallery Modality   #Description   #Image
-Market-1501 zheng2015scalable Image (Cropped) Image (Cropped) 32,668
-CUHK03 li2014deepreid Image (Cropped) Image (Cropped) 14,097
-CUHK-SYSU xiao2017joint Image (Cropped) Image (Full) 18,184
-PRW zheng2017person Image (Cropped) Image (Full) 11,816
-CUHK-PEDES li2017person Text (Appearance) Image (Cropped) 80,412 40,206
-ICFG-PEDES ding2021semantically Text (Appearance) Image (Cropped) 54,522 54,522
-CUHK-SYSU-TBPS zhang2023text Text (Appearance) Image (Full) 35,960 18,184
-PRW-TBPS zhang2023text Text (Appearance) Image (Full) 19,009 11,246
+  Market-1501 [zheng2015scalable]   Image (Cropped)   Image (Cropped)   –   32,668
+  CUHK03 [li2014deepreid]   Image (Cropped)   Image (Cropped)   –   14,097
+  CUHK-SYSU [xiao2017joint]   Image (Cropped)   Image (Full)   –   18,184
+  PRW [zheng2017person]   Image (Cropped)   Image (Full)   –   11,816
+  CUHK-PEDES [li2017person]   Text (Appearance)   Image (Cropped)   80,412   40,206
+  ICFG-PEDES [ding2021semantically]   Text (Appearance)   Image (Cropped)   54,522   54,522
+  CUHK-SYSU-TBPS [zhang2023text]   Text (Appearance)   Image (Full)   35,960   18,184
+  PRW-TBPS [zhang2023text]   Text (Appearance)   Image (Full)   19,009   11,246
   ScenePerson-13W   Text (Appearance+scene)   Image (Full)   131,438   131,438
-2 ScenePerson-13W Dataset
-The ScenePerson-13W dataset is proposed to support scene-aware person retrieval by incorporating rich scene context alongside detailed pedestrian appearance in textual descriptions.
+IV ScenePerson-13W Dataset
+The ScenePerson-13W dataset is proposed to support scene-aware text-based person retrieval by incorporating rich scene context alongside detailed pedestrian appearance in textual descriptions.
 It contains $131,438$ pedestrian instances across $112,872$ diverse scenes, with each instance accompanied by a textual description, a full-scene image, and a bounding box highlighting the pedestrian’s location.
-The dataset is split into a training set of $121,438$ images and a testing set of $10,000$ images, which preserves the complex scene structures and identity distributions while reducing redundancy and ensures comprehensive evaluation across diverse scenarios.
+The dataset consists of $121,438$ training images and $10,000$ test images, which preserves scene complexity and pedestrian distributions with reduced redundancy, enabling comprehensive evaluation across diverse scenarios.
 The construction process of the dataset illustrated in Figure 3, involves five key steps: scene segmentation, pedestrian detection and tracking, completeness filtering, deduplication in image collection, and textual description generation.
-More examples from the dataset are presented in Appendix A.3.
 Image Collection.
-The image data in ScenePerson-13W is constructed from the QVHighlights video dataset lei2107qvhighlights through a multi-stage processing pipeline designed to ensure data quality and diversity.
-First, videos are segmented into distinct scenes, each representing a unique environmental context.
+The image data in ScenePerson-13W is constructed from the QVHighlights video dataset [lei2107qvhighlights] through a multi-stage processing pipeline designed to ensure data quality and diversity.
+1) Scene segmentation.
+Videos are segmented into distinct scenes, each representing a unique environmental context.
 This segmentation enables the dataset to capture a wide range of pedestrian-environment combinations, supporting comprehensive scene-aware analysis.
-Second, the pedestrian detection and tracking are performed using YOLOv11 khanam2024yolov11 .
+2) Pedestrian detection and tracking.
+Each segmented scene is then processed to detect and track pedestrians.
 For each detected pedestrian instance, the frame with the highest confidence score is selected, and the corresponding full scene image with cropped pedestrian region are retained.
-Third, to enhance data quality, incomplete or noisy detections–such as those capturing only partial body parts–are filtered out using MediaPipe keypoint detection mediapipe_pose .
+3) Completeness filtering.
+To enhance data quality, incomplete or noisy detections–such as those capturing only partial body parts–are filtered out using MediaPipe keypoint detection [mediapipe_pose].
 Only images containing a head and at least one shoulder are preserved, ensuring the availability of essential visual features.
+4) Deduplication.
 Finally, a deduplication process is applied to reduce redundancy.
-Using CLIP radford2021learning , visual similarity between images is assessed.
+Using CLIP [radford2021learning], visual similarity between images is assessed.
 Instances that are highly similar in both the cropped pedestrian region and the full scene are discarded, promoting the uniqueness and diversity of the resulting dataset.
 Finally, we collect $131,438$ pedestrian instance, accompanied with $131,438$ corresponding scene images.
+Figure 4: Sunburst visualization of ScenePerson-13W, showing hierarchical feature distribution across Appearance, Landmarks, and Proximity.
 Description Generation.
-For each of the $131,438$ retained pedestrian instances, a corresponding textual description is generated using InternVL2.5-8B chen2024expanding .
+For each of the $131,438$ retained pedestrian instances, a corresponding textual description is generated using InternVL2.5-8B [chen2024expanding].
 Specifically, the model is prompted with the full-scene image in which the target pedestrian is highlighted using a red bounding box.
 The prompt is structured as follows:
-Generate a description of the individual within the red bounding box and the connection with the surroundings, in a style similar to the template: {template}.
-Ensure that the description is concise within $50$ words and accurate.
+Generate a description of the individual within the red bounding box and the connection with the surroundings.
+Ensure that the description is a natural, fluent paragraph (under $50$ words) that a real witness would actually say to police or friends when trying to help find this exact person.
 Do not include text details or labels, and strictly avoid any punctuation other than commas and periods.
 The prompt is designed to elicit detailed descriptions by guiding the model to focus on three key aspects: the individual’s appearance, location within the scene, and relationships with surrounding elements, following a structured template.
-Compared to existing person retrieval datasets, ScenePerson-13W offers a significant advancement by integrating fine-grained appearance details with rich global scene context across diverse scenarios, as shown in Table  1.
+Compared to existing person retrieval datasets[zheng2015scalable, li2014deepreid, xiao2017joint, zheng2017person, li2017person, ding2021semantically, zhang2023text], ScenePerson-13W offers a significant advancement by integrating fine-grained appearance details with rich global scene context across diverse scenarios and surpassing prior datasets in scale, as shown in Table  II.
+And figure 4 further illustrates the hierarchical feature distribution in ScenePerson-13W, enabling precise, context-rich descriptions that jointly capture fine-grained appearance and complex scene context in real-world settings.
 Figure 5: Overview of the proposed SA-Person framework.
@@
 The second stage, scene-aware reranking, re-ranks the top-$K$ candidates using scene-aided text and the full-scene image.
-3 Method
-To tackle the challenges of holistic understanding in both pedestrian appearance and scene context, we propose SA-Person, a two-stage framework, as shown in Figure 2.
-The first stage aligns pedestrian-specific regions with appearance-related descriptions to retrieve candidates.
-The second stage refines the initial candidates by incorporating global scene context.
-3.1 Appearance-based Retrieval
+V Method
+Motivated by the insights revealed in Section III, we propose SA-Person, a straightforward yet effective two-stage framework, as shown in Figure 5.
+The first stage aligns pedestrian-specific regions with appearance-related textual descriptions to retrieve a compact set of candidates.
+The second stage refines the top candidates by incorporating global scene context via a training-free scene-aware re-ranking method.
+V-A Appearance-based Retrieval
 Appearance Embedding Training.
 The first stage focuses on retrieving a shortlist of candidate individuals from the gallery set based solely on their visual appearance, given the appearance-related content of the textual query.
-For a gallery of full-scene images $\mathcal{G}=\{I_{1},I_{2},\dots,I_{N}\}$ and a textual query $T$ describing the target person, we begin by extracting the appearance-related content from $T$.
-Specifically, we parse each query description in the dataset and extract the segments of the description that describes the target person’s visual traits, such as clothing, hairstyle, or accessories (e.g., “A man wearing a red jacket and blue jeans is laughing”), denoted as ${T}_{\text{app}}$.
+For a gallery of full-scene images $\mathcal{G}=\{I_{1},I_{2},\dots,I_{N}\}$ and a text $T$ describing the target person, we use appearance-related text that describes the target person’s visual traits, such as clothing, hairstyle, or accessories (e.g., A man wearing a red jacket and blue jeans is laughing), denoted as ${T}_{\text{app}}$.
 This ensures that the retrieval process prioritizes appearance-focused matching, excluding contextual or environmental cues at this stage.
@@
 where $\text{Enc}_{\text{v}}$ and $\text{Enc}_{\text{t}}$ represent visual encoder and textual encoder, respectively, $f_{ij}^{v}\in\mathbb{R}^{d}$ represents the visual feature representation of the $j$-th person in image $I_{i}$ and $f^{t}\in\mathbb{R}^{d}$ is the textual feature representation aligned with the visual feature space.
-To enhance cross-modal alignment, we fine-tune the dual-encoder on these appearance-related image-text pairs from ScenePerson-13W, following the implicit relation reasoning proposed in jiang2023cross .
+To enhance cross-modal alignment, we fine-tune the dual-encoder on these appearance-related image-text pairs from ScenePerson-13W, following the implicit relation reasoning proposed in [jiang2023cross].
 The loss is formally expressed as:
-$\mathcal{L}_{\text{app}}=\mathbb{E}_{(b_{ij},\mathcal{T}_{\text{app}},y_{ij})\sim\mathcal{D}_{\text{app}}}\left[\ell_{\text{IRR}}(f_{ij}^{v},f^{t},y_{ij};\Theta)+\ell_{\text{SDM}}(f_{ij}^{v},f^{t},y_{ij};\Theta)+\ell_{\text{ID}}(f_{ij}^{v},f^{t},y_{ij};\Theta)\right],$ (2)
-where $\mathcal{D}_{\text{app}}$ is the set of appearance-related image-text pairs with identity labels $y_{ij}$, $\ell_{\text{IRR}}$, $\ell_{\text{SDM}}$, and $\ell_{\text{ID}}$ are the respective loss components as defined in  jiang2023cross , and $\Theta$ denotes the parameters of the image and text encoders.
+$\begin{split}\mathcal{L}_{\text{app}}=\mathbb{E}_{(b_{ij},\mathcal{T}_{\text{app}},y_{ij})\sim\mathcal{D}_{\text{app}}}\big[&\ell_{\text{IRR}}(f_{ij}^{v},f^{t},y_{ij};\Theta)\\ +&\ell_{\text{SDM}}(f_{ij}^{v},f^{t},y_{ij};\Theta)\\ +&\ell_{\text{ID}}(f_{ij}^{v},f^{t},y_{ij};\Theta)\big],\end{split}$ (2)
+where $\mathcal{D}_{\text{app}}$ is the set of appearance-related image-text pairs with identity labels $y_{ij}$, $\ell_{\text{IRR}}$, $\ell_{\text{SDM}}$, and $\ell_{\text{ID}}$ are the respective loss components as defined in [jiang2023cross], and $\Theta$ denotes the parameters of the image and text encoders.
 The expectation is computed over the training set, ensuring discriminative and robust feature representations.
 Retrieval.
-During inference, we first apply a YOLO-based person detector to a gallery of full-scene images $\mathcal{G}=\{\mathbf{I}_{j}\}_{j=1}^{M}$, extracting all detected person crops across the gallery.
+During inference, we first apply the pedestrian detector to a gallery of full-scene images $\mathcal{G}=\{\mathbf{I}_{j}\}_{j=1}^{M}$, extracting all detected person crops across the gallery.
 This yields a set of $K$ cropped person images $\mathcal{C}=\{\mathbf{C}_{k}\}_{k=1}^{K}$, where each $\mathbf{C}_{k}$ originates from a full-scene image $\mathbf{I}_{j}\in\mathcal{G}$.
 For each crop $\mathbf{C}_{k}$, we also retain its bounding box $\mathbf{B}_{k}$ and its source image identifier $j=s(k)$, indicating that $\mathbf{C}_{k}$ was cropped from $\mathbf{I}_{s(k)}$.
-Each cropped image $C_{k}$ is encoded into $f_{k}^{v}$ using the fine-tuned image encoder and the appearance-related textual query $\mathcal{T}_{\text{app}}$ is encoded into $f^{t}$, as defined previously in Eq.(1).
+Each cropped image $C_{k}$ is encoded into $f_{k}^{v}$ using the fine-tuned image encoder and the appearance-related text $\mathcal{T}_{\text{app}}$ is encoded into $f^{t}$, as defined previously in Eq.(1).
 The similarity score between the textual feature and each person image is computed as:
@@
 $\mathcal{C}_{\text{top}}$ will be passed to the subquent re-ranking stage for further improvement based on scene understanding.
-3.2 Scene-aware Re-ranking
-While the appearance-based retrieval stage provides a strong initial filter by matching at the region-level focused on appearance, it lacks the ability to incorporate broader contextual cues, such as spatial configurations, interactions, or referential expressions.
-To complement this, we introduce a scene-aware ranking module based on InternVL 2.5-8B chen2024expanding , an advanced multimodal large language model (MLLM) capable of joint visual-linguistic reasoning over complex scenes and full descriptions, without any task-specific fine-tuning.
+V-B Scene-aware Re-ranking
+While the appearance-based retrieval stage provides a strong initial filter by matching at the region-level focused on appearance, it can be limited in fully leveraging broader contextual cues compared to MLLMs, which offer superior holistic reasoning in full-scene settings.
+To leverage these complementary strengths without the overhead of gallery-wide processing, we introduce SceneRanker, a scene-aware ranking module that applies a MLLM to refine the top candidates from the appearance-based retrieval stage, without any task-specific fine-tuning.
+This confines expensive inference to a compact subset for scalable disambiguation.
 Candidate-to-Scene Mapping.
@@
 Scene-aware Multimodal Prompt.
-To guide SceneRanker to perform reranking, we construct a unified prompt that integrates both the annotated visual inputs and an instructional textual query.
-The MLLM input is a structured multimodal prompt combining the $K$ full-scene images $\{I_{m}\}_{m=1}^{K}$, their corresponding bounding box annotations $\{\mathbf{B}_{m}\}_{m=1}^{K}$, and the complete textual query $T$.
+To guide SceneRanker to perform reranking, we construct a unified prompt that integrates both the annotated visual inputs and an instructional text.
+The MLLM input is a structured multimodal prompt combining the $K$ full-scene images $\{I_{m}\}_{m=1}^{K}$, their corresponding bounding box annotations $\{\mathbf{B}_{m}\}_{m=1}^{K}$, and the complete text $T$.
 The visual prompt is formatted as a markdown sequence, where each full-scene image is paired with its bounding box annotation:
 $\mathbf{P}_{v}=\text{Concat}\left(\{\text{Image-}m:\langle\text{image}\rangle\langle\text{box}\rangle\mathbf{B}_{m}\langle/\text{box}\rangle\}_{m=1}^{K}\right),$
-[…]
-Image-1: <image><box>$B_{1}$</box> Image-2: <image><box>$B_{2}$</box> Image-$K$: <image><box>$B_{K}$</box> For the text description: ‘$T$’, order the images based on how accurately they reflect overall context in the text, starting with the most faithful match.
+where $\langle\text{image}\rangle$ is a placeholder for $\mathbf{I}_{m}$, and $\langle\text{box}\rangle\mathbf{B}_{m}\langle/\text{box}\rangle$ denotes the spatial region of interest corresponding to the candidate.
+And an instructional prompt is appended to define the ranking task.
+Specifically, the final multimodal prompt $\mathbf{P}$ takes the following form:
+Image-1: <image> $B_{1}</box$> Image-2: <image> $B_{2}</box$> Image-3: <image> $B_{3}</box$> Image-K: <image> $B_{K}</box$> Instruction: For the text description: “TT”, order the images based on how accurately they reflect the overall context in the text, starting with the most faithful match.
 Please output only a Python list format like this: [4, 2, 8, 1, 5, 3, 6, 9, 10, 7, $\cdots$, $K$].
-Re-ranking. The SceneRanker processes the multimodal input ($(\{\mathbf{I}_{m},\mathbf{B}_{m}\}_{m=1}^{K},\mathbf{P})$ to produce a ranked list of indices:
+Re-ranking. The SceneRanker processes the multimodal input $(\{\mathbf{I}_{m},\mathbf{B}_{m}\}_{m=1}^{K},\mathbf{P})$ to produce a ranked list of indices:
 $\mathbf{R}=\text{SceneRanker}(\{\mathbf{I}_{m},\mathbf{B}_{m}\}_{m=1}^{K},\mathbf{P}),$
@@
 The reranking output serves as the final retrieval results of our framework, integrating both localized appearance matching and global scene understanding.
-4 Experiments
-4.1 Experimental Setup
+VI Experiments
+VI-A Experimental Setup
 Dataset and Evaluation Metrics.
-All experiments are conducted on the proposed ScenePerson-13W dataset, which is divided into $121,438$ image-text pairs for training and $10,000$ image-text pairs for testing.
-Following prior retrieval works bai2023rasa ; cao2024empirical , we evaluate overall retrieval performance using Rank@k (R@k) for k=1, 5, and 10, and mean Average Precision (mAP).
+Experiments are conducted on four benchmarks: the proposed ScenePerson-13W dataset, augmented CUHK-SYSU [xiao2017joint], Urban1K [zhang2024long], and a 5K subset of ShareGPT4V [chen2024sharegpt4v].
+Unless otherwise specified, performance metrics are reported on ScenePerson-13W.
+Following prior retrieval works [bai2023rasa, cao2024empirical], we evaluate overall retrieval performance using Rank@k (R@k) for k=1, 5, and 10, respectively, and mean Average Precision (mAP).
 R@k measures the proportion of queries where the target individual is correctly identified within the top k candidates, while mAP quantifies the ranking quality across all queries.
+TABLE III: Comparison of the proposed SA-Person with state-of-the-art methods.
+Method Venue Focus Backbone R@1 (%) R@5 (%) R@10 (%) mAP
+Region-level:
+RaSa [bai2023rasa] IJCAI’23 Appearance ViT-B/16 62.51 83.83 88.09 72.05
+IRRA [jiang2023cross] CVPR’23 Appearance ViT-B/16 60.57 83.35 87.86 70.65
+IRRA [jiang2023cross] CVPR’23 Appearance ViT-L/14 63.37 84.54 88.67 72.71
+TBPS-CLIP [cao2024empirical] AAAI’24 Appearance ViT-B/16 50.29 75.24 81.58 61.54
+RDE [qin2024noisy] CVPR’24 Appearance ViTB/16 61.26 83.50 88.23 71.17
+MARS [ergasti2024mars] TOMM’24 Appearance ViTB/16 62.93 84.26 88.65 72.43
+FMFA [yin2025cross] arXiv’25 Appearance ViTB/16 59.77 83.22 88.05 70.09
+Global-level:
+CLIP [radford2021learning] ICML’21 Scene-aided ViT-B/16 56.37 80.73 87.16 67.22
+CLIP [radford2021learning] ICML’21 Scene-aided ViT-L/14 63.51 85.53 90.28 73.33
+LongCLIP [zhang2024long] ECCV’24 Scene-aided ViT-B/16 58.90 82.41 88.23 69.42
+LongCLIP [zhang2024long] ECCV’24 Scene-aided ViT-L/14 65.75 86.22 90.95 75.05
+Region-to-Global Cascade:
+SA-Person Ours Scene-aided ViT-B/16 77.42 86.61 87.86 81.96
+SA-Person Ours Scene-aided ViT-L/14 78.34 87.54 88.67 82.86
 Implementation Details.
-For the appearance-based retrieval stage, we adopt YOLOv5wu2021application for pedestrian detection and MediaPipe mediapipe_pose for filtering to ensure valid person instances.
-The dual-encoder model is fine-tuned following the original IRRA jiang2023cross implementation, initialized with CLIP-ViT-B/16 radford2021learning for images and CLIP text Transformer radford2021learning for text.
-The model is trained for $60$ epochs on the ScenePerson-13W dataset using the Adam optimizer with an initial learning rate of $10^{-5}$ and a cosine learning rate decay schedule.
+For the appearance-based retrieval stage, the pedestrian detector is composed of YOLOv5 [wu2021application] for initial detection and MediaPipe [mediapipe_pose] for keypoint filtering, ensuring valid and reliable person instances.
+The dual-encoder model is fine-tuned following the original IRRA [jiang2023cross] implementation, initialized with CLIP-ViT-B/16 [radford2021learning] for images and CLIP text Transformer [radford2021learning] for text.
+The model is trained for $60$ epochs using the Adam optimizer with an initial learning rate of $10^{-5}$ and a cosine learning rate decay schedule.
 The image resolution of input crops is resized to $384\times 128$ pixels.
 During inference, the size of candidate set $K$ selected for further processing is set to $10$.
-For the SceneRanker, we use InternVL 2.5-8B chen2024expanding without task-specific fine-tuning.
-The generation is performed with a maximum output length of $128$ tokens and deterministic decoding.
-The maximum number of patches that an input image is divided into during preprocessing is set to $10$ to balance computational efficiency and detailed visual analysis.
-All other generation hyperparameters retain their default values as defined by the Hugging Face Transformers wolf2019huggingface library.
+For the scene-aware re-ranking stage, we employ InternVL 2.5-8B [chen2024expanding] as SceneRanker, without task-specific fine-tuning.
+Inference is performed with a maximum of $128$ output tokens and deterministic decoding.
+All other settings follow the default Hugging Face Transformers [wolf2019huggingface] configuration.
 All experiments are conducted on four RTX A40 GPUs with 48 GB of memory each.
-4.2 Comparison with State-of-the-art Methods
-We compare the proposed SA-Person method with a range of state-of-the-art retrieval baselines, grouped by their retrieval strategies and visual grounding scope.
+VI-B Comparison with State-of-the-art Methods
+We compare the proposed SA-Person with a range of state-of-the-art retrieval baselines, grouped by their retrieval strategies and visual grounding scope.
 Region-level methods solely focus on cropped pedestrian images, emphasizing appearance cues but lacking broader scene context.
@@
 In contrast, SA-Person adopts a region-to-global retrieval paradigm, initially matching pedestrian-specific regions to ensure discriminative grounding, followed by scene-aware reranking to incorporate semantic information from the broader environment.
-As shown in Table 2, SA-Person consistently outperforms all baselines by a large margin.
-For region-level methods, it yields a significant boost of 15.41% at R@1 and 10.43% at mAP.
-Against the strongest region-level model, IRRA (ViT-L/14), SA-Person still improves R@1 by 14.97% and mAP by 10.15%.
+As shown in Table III, SA-Person consistently outperforms all baselines by a large margin.
+For region-level methods, SA-Person yields a significant boost of 14.97% at R@1 and 10.15% at mAP over the strongest method IRRA (ViT-L/14).
 For global-level methods, SA-Person exceeds CLIP (ViT-L/14) by 14.83% at R@1 and 9.53 at mAP, and outperforms the best-performing LongCLIP (ViT-L/14) by 12.59% at R@1 and 7.81% at mAP.
 These results collectively highlight the superiority of our region-to-global cascade retrieval strategy, which integrates region-level discrimination with scene-level context and enhance both localization and semantic reasoning for more robust text-to-person retrieval.
-Table 2: Comparison of the proposed SA-Person with state-of-the-art methods.
-Method Focus Backbone R@1 (%) R@5 (%) R@10 (%) mAP
-Region-level:
-RaSa bai2023rasa Appearance ViT-B/16 62.51 83.83 88.09 72.05
-IRRA jiang2023cross Appearance ViT-B/16 60.57 83.35 87.86 70.65
-IRRA jiang2023cross Appearance ViT-L/14 63.37 84.54 88.67 72.71
-TBPS-CLIP cao2024empirical Appearance ViT-B/16 50.29 75.24 81.58 61.54
-RDE qin2024noisy Appearance ViTB/16 61.26 83.50 88.23 71.17
-MARS ergasti2024mars Appearance ViTB/16 62.93 84.26 88.65 72.43
-Global-level:
-CLIP radford2021learning Scene-aided ViT-B/16 56.37 80.73 87.16 67.22
-CLIP radford2021learning Scene-aided ViT-L/14 63.51 85.53 90.28 73.33
-LongCLIP zhang2024long Scene-aided ViT-B/16 58.90 82.41 88.23 69.42
-LongCLIP zhang2024long Scene-aided ViT-L/14 65.75 86.22 90.95 75.05
-Region-to-Global Cascade:
-SA-Person (Ours) Scene-aided ViT-B/16 77.42 86.61 87.86 81.96
-SA-Person (Ours) Scene-aided ViT-L/14 78.34 87.54 88.67 82.86
-Table 3: Performance comparison of text-based person retrieval with and without the proposed SceneRanker. Red numbers indicate the absolute improvement achieved by our SceneRanker.
+TABLE IV: Performance comparison of text-based person retrieval methods with and without the proposed SceneRanker. Red numbers indicate the absolute improvement achieved by our SceneRanker.
 Methods R@1 R@5 R@10 mAP
-     TBPS-CLIP cao2024empirical      50.29      75.24      81.58      61.54
+TBPS-CLIP [cao2024empirical] 50.29 75.24 81.58 61.54
 +SceneRanker $72.25_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+21.96}}$ $80.23_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+4.99}}$ $81.58_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+0.00}}$ $76.54_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+15.00}}$
-     IRRA(ViT-B/16) jiang2023cross      60.57      83.35      87.86      70.65
+IRRA(ViT-B/16) [jiang2023cross] 60.57 83.35 87.86 70.65
 +SceneRanker $77.42_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+16.85}}$ $86.61_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+3.26}}$ $87.86_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+0.00}}$ $81.96_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+11.31}}$
-     IRRA(ViT-L/14) jiang2023cross      63.37      84.54      88.67      72.71
+IRRA(ViT-L/14) [jiang2023cross] 63.37 84.54 88.67 72.71
 +SceneRanker $\textbf{78.34}_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+14.97}}$ $\textbf{87.54}_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+3.00}}$ $\textbf{88.67}_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+0.00}}$ $\textbf{82.86}_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+10.15}}$
-     RDE qin2024noisy      61.26      83.50      88.23      71.17
-      +SceneRanker      $77.58_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+16.32}}$      $86.84_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+3.34}}$      $88.23_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+0.00}}$      $82.18_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+11.01}}$
-     RaSa bai2023rasa      62.51      83.83      88.09      72.05
-      +SceneRanker      $77.84_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+15.33}}$      $86.83_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+3.00}}$      $88.09_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+0.00}}$      $82.27_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+10.22}}$
-     MARS ergasti2024mars      62.93      84.26      88.65      72.43
-      +SceneRanker      $78.32_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+15.39}}$      $87.39_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+3.13}}$      $88.65_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+0.00}}$      $82.77_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+10.34}}$
-4.3 Compatibility of the Proposed SceneRanker
-The proposed SceneRanker is highly compatible and can be seamlessly adapted to various text-based person retrieval methods.
+RDE [qin2024noisy] 61.26 83.50 88.23 71.17
++SceneRanker $77.58_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+16.32}}$ $6.84_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+3.34}} 88$$8.23_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+0.00}}$ $82.18_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+11.01}}$
+RaSa [bai2023rasa] 62.51 83.83 88.09 72.05
++SceneRanker $77.84_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+15.33}}$ $6.83_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+3.00}} 88$$8.09_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+0.00}}$ $82.27_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+10.22}}$
+MARS [ergasti2024mars] 62.93 84.26 88.65 72.43
++SceneRanker $78.32_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+15.39}}$ $7.39_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+3.13}} 88$$8.65_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+0.00}}$ $82.77_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+10.34}}$
+FMFA [yin2025cross] 59.77 83.22 88.05 70.09
++SceneRanker $77.35_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+17.58}}$ $6.59_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+3.37}} 88$$8.05_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+0.00}}$ $81.95_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+11.86}}$
+VI-C Compatibility of the Proposed SceneRanker
+The proposed SceneRanker, exploiting global scene context for precise disambiguation in the scene-aware reranking stage, is highly compatible and can be seamlessly adapted to various text-based person retrieval methods.
 Specifically, these methods can function as the first stage of appearance-based retrieval in the SA-Person framework.
-We evaluate the compatibility of SceneRanker in Table 3, where it consistently demonstrates performance improvements.
-Across all baselines, SceneRanker delivers over a 15% improvement in R@1 and more than a 10% increase in mAP.
-Notably, it achieves significant gains for TBPS-CLIP, with R@1 increasing by 21.93% (from 50.32% to 72.25%).
+We evaluate the compatibility of SceneRanker in Table IV, where it consistently demonstrates performance improvements.
+Across all baselines, SceneRanker delivers over a 14% improvement in R@1 and more than a 10% increase in mAP.
+Notably, it achieves significant gains for TBPS-CLIP, with R@1 increasing by 21.96% (from 50.29% to 72.25%).
 These consistent enhancements highlight SceneRanker as a flexible and effective component that integrates seamlessly with diverse text-based person retrieval methods.
 By leveraging complementary contextual cues beyond standalone pedestrian appearance, SceneRanker refines candidate selection and significantly improves retrieval results.
-4.4 Ablation Studies
-We conduct ablation studies to evaluate the effectiveness of the proposed SceneRanker, focusing on (1) the prompt design strategy, (2) the choice of multimodal large language models (MLLMs), and (3) the candidate size in the first-stage retrieval.
-Table 4: Comparison the impact of using different MLLMs into the SceneRanker.
-Due to varying MLLM capabilities in processing structured prompts, we adopt Naive Prompt across all models to ensure fair comparison.
+VI-D Ablation Studies
+We conduct ablation studies to evaluate the effectiveness of the proposed SceneRanker, focusing on (1) the prompt design strategy, (2) the choice of MLLMs, and (3) the candidate size in the first-stage retrieval.
+Different Prompt Design. We compare three strategies designed to enhance spatial grounding of pedestrian-specific region information in scene-aware multimodal prompts for person retrieval:
+TABLE V: Performance of SceneRanker with various text-based person retrieval methods under different prompt designs. Red numbers indicate the improvement brought by each prompt design.
+Methods R@1 R@5 R@10 mAP
+TBPS-CLIP [cao2024empirical] 50.29 75.24 81.58 61.54
++NP $71.41_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+21.12}}$ $80.19_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+4.95}}$ $81.58_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+0.00}}$ $76.03_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+14.49}}$
++BOP $70.77_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+20.48}}$ $80.16_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+4.92}}$ $81.58_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+0.00}}$ $75.69_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+14.15}}$
++BEP $\textbf{72.25}_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+21.96}}$ $\textbf{80.23}_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+4.99}}$ $\textbf{81.58}_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+0.00}}$ $\textbf{76.54}_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+15.00}}$
+IRRA(ViT-B/16) [jiang2023cross] 60.57 83.35 87.86 70.65
++NP $76.80_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+16.23}}$ $86.41_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+3.06}}$ $87.86_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+0.00}}$ $81.53_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+10.88}}$
++BOP $75.78_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+15.21}}$ $86.39_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+3.04}}$ $87.86_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+0.00}}$ $80.99_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+10.34}}$
++BEP $\textbf{77.42}_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+16.85}}$ $\textbf{86.61}_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+3.26}}$ $\textbf{87.86}_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+0.00}}$ $\textbf{81.96}_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+11.31}}$
+IRRA(ViT-L/14) [jiang2023cross] 63.37 84.54 88.67 72.71
++NP ${77.55}_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+14.18}}$ ${87.42}_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+2.88}}$ ${88.67}_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+0.00}}$ ${82.33}_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+9.62}}$
++BOP ${76.69}_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+13.32}}$ ${87.29}_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+2.75}}$ ${88.67}_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+0.00}}$ ${81.85}_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+9.14}}$
++BEP $\textbf{78.34}_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+14.97}}$ $\textbf{87.54}_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+3.00}}$ $\textbf{88.67}_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+0.00}}$ $\textbf{82.86}_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+10.15}}$
+RDE [qin2024noisy] 61.26 83.50 88.23 71.17
++NP $77.13_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+15.87}}$ $6.83_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+3.33}} 88$$8.23_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+0.00}}$ $81.87_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+10.70}}$
++BOP $76.29_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+15.03}}$ $\textbf{86.88}_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+3.38}}$ $88.23_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+0.00}}$ $81.46_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+10.29}}$
++BEP $\textbf{77.58}_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+16.32}}$ ${86.84}_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+3.34}}$ $\textbf{88.23}_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+0.00}}$ $\textbf{82.18}_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+11.01}}$
+RaSa [bai2023rasa] 62.51 83.83 88.09 72.05
++NP $77.04_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+14.53}}$ $6.66_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+2.83}} 88$$8.09_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+0.00}}$ $81.76_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+9.71}}$
++BOP $76.24_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+13.73}}$ $6.74_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+2.91}} 88$$8.09_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+0.00}}$ $81.29_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+9.24}}$
++BEP $\textbf{77.84}_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+15.33}}$ $\textbf{86.83}_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+3.00}}$ $\textbf{88.09}_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+0.00}}$ $\textbf{82.27}_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+10.22}}$
+MARS [ergasti2024mars] 62.93 84.26 88.65 72.43
++NP $77.26_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+14.33}}$ $7.30_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+3.04}} 88$$8.65_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+0.00}}$ $82.10_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+9.67}}$
++BOP $76.57_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+13.64}}$ $7.14_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+2.88}} 88$$8.65_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+0.00}}$ $81.69_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+9.26}}$
++BEP $\textbf{78.32}_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+15.39}}$ $\textbf{87.39}_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+3.13}}$ $\textbf{88.65}_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+0.00}}$ $\textbf{82.77}_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+10.34}}$
+FMFA [yin2025cross] 59.77 83.22 88.05 70.09
++NP $76.51_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+16.74}}$ $6.58_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+3.36}} 88$$8.05_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+0.00}}$ $81.47_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+11.38}}$
++BOP $75.83_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+16.06}}$ $6.48_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+3.26}} 88$$8.05_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+0.00}}$ $81.09_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+11.00}}$
++BEP $\mathbf{77.35}_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+17.58}}$ $\mathbf{86.59}_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+3.37}}$ $\mathbf{88.05}_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+0.00}}$ $\mathbf{81.95}_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+11.86}}$
+1) Naive Prompt (NP), which provides only the textual description and the full-scene image without explicit region information; 2) Box Overlay Prompt (BOP), which visually highlights the candidate region with a red rectangle on the full-scene image; 3) Box Embedded Prompt (BEP), which includes the bounding box coordinates of the candidate region in the multimodal prompt using <box> tags.
+As shown in Table V, SceneRanker significantly boosts performance for all methods, with R@1 improvements exceeding above 14.97% with BEP, underscoring its compatibility and effectiveness in refining appearance-based retrieval.
+Across all methods, BEP consistently outperforms NP and BOP, achieving the highest R@1, suggesting that embedding bounding box coordinates enhances spatial grounding, enabling better alignment with textual descriptions.
+Notably, BOP even play worse than NP across all baselines, with a representative drop of 1.02% at R@1 and 0.51% at mAP on IRRA with ViT-B/16, indicating that direct visual highlighting may introduce visual distractions that disrupt holistic scene understanding.
+Different Multimodal Large Language Models.
+We evaluate the impact of integrating different 8B-scale MLLMs into the proposed SceneRanker, using the unified naive prompt to mitigate varying capabilities in processing structured inputs.
+As shown in Table VI, all three models, Qwen-VL 2.5, MiniCPM-V, and InternVL 2.5, consistently improve over the IRRA baseline, with respective R@1 gains of +11.7%, +7.96%, and +14.18%.
+In addition, we observe a strong correlation between model capacity and performance.
+Taking InternVL2.5 as an example, scaling from 4B to 8B brings further improvement, with R@1 increasing from 75.21% to 77.55%.
+In contrast, smaller models below 4B such as InternVL 1B and 2B show significantly lower performance, indicating insufficient multimodal understanding.
+These results confirm SceneRanker’s effectiveness and generalizability across MLLM architectures with sufficient vision-language reasoning capacity.
+TABLE VI: Impact of integrating different MLLMs into SceneRanker. Due to varying MLLM capabilities in processing structured prompts, we adopt Naive Prompt across all models to ensure fair comparison.
 Model Parameters R@1 R@5 R@10 mAP
-IRRA (ViT-L/14) 63.37 84.54 88.67 72.71
-QwenVL2.5 bai2025qwen2 3 Billion 54.57 78.23 88.67 65.04
-8 Billion 75.07 86.34 88.67 80.45
-MiniCPM-V-2_6 yao2024minicpm 8 Billion 71.33 84.66 88.67 77.78
-InternVL2.5 chen2024expanding 1 Billion 3.01 80.14 88.67 26.71
+IRRA (ViT-L/14) 590 Million 63.37 84.54 88.67 72.71
+QwenVL2.5 [bai2025qwen2] 3 Billion 54.57 78.23 88.67 65.04
+QwenVL2.5 [bai2025qwen2] 8 Billion 75.07 86.34 88.67 80.45
+MiniCPM-V-2_6 [yao2024minicpm] 8 Billion 71.33 84.66 88.67 77.78
+InternVL2.5 [chen2024expanding] 1 Billion 3.01 80.14 88.67 26.71
 2 Billion 44.14 75.32 88.67 56.74
@@
 8 Billion 77.55 87.42 88.67 82.33
-Different Prompt Design. We compare three strategies designed to enhance spatial grounding of pedestrian-specific region information in scene-aware multimodal prompts for person retrieval:
-Prompt Design R@1 R@5 R@10 mAP 63.37 84.54 88.67 72.71 Naive Prompt 77.55 87.42 88.67 82.33 Box Overlay Prompt 76.69 87.29 88.67 81.85 Box Embedded Prompt 78.34 87.54 88.67 82.86
-Table 5: Impact of visual prompt design on the performance of SceneRanker. IRRA(ViT-L/14) is adopted as the baseline for comparison, without prompt design.
-1) Naive Prompt (NP), which provides only the textual description and the full-scene image without explicit region information; 2) Box Overlay Prompt (BOP), which visually highlights the candidate region with a red rectangle on the full-scene image; 3) Box Embedded Prompt (BEP), which includes the bounding box coordinates of the candidate region in the multimodal prompt using <box> tags.
-The complete designs of the three prompt strategies are provided in Appendix A.5.
-As shown in Table 5, incorporating box embedded prompt into the proposed SceneRanker outperforms naive prompt by 0.79% at R@1 and 0.53% at mAP, confirming that explicitly encoding region locations strengthens the model’s spatial grounding.
-Compared to the box overlay prompt, box embedded prompt achieves an even greater improvement of 1.65% at R@1 and 1.01% at mAP, suggesting that structured coordinate inputs are more effective than visual overlays.
-Notably, BOP even play worse than NP by 0.86% at R@1 and 0.48% at mAP, indicating that direct visual highlighting may introduce visual distractions that interfere with the model’s scene understanding.
-Different Multimodal Large Language Models.
-We evaluate the effect of integrating different 8B-scale MLLMs into the proposed SceneRanker, using the same naive prompt for fair comparison.
-As shown in Table 4, all three models, Qwen-VL 2.5, MiniCPM-V, and InternVL 2.5, consistently improve over the IRRA baseline, with respective R@1 gains of 11.7%, +7.96%, and +14.18%.
-In addition, we observe a strong correlation between model capacity and performance.
-Taking InternVL 2.5 as an example, scaling from 4B to 8B brings further improvement, with R@1 increasing from 75.21% to 77.55%.
-In contrast, smaller models below 4B such as InternVL 1B and 2B show significantly lower performance, indicating insufficient multimodal understanding.
-These findings suggest that various MLLMs are well-suited for SceneRanker, and its performance benefits from stronger vision-language reasoning capabilities.
-Importantly, this holds true regardless of the specific architecture, provided the model has sufficient capacity.
+Figure 6: Impact of candidate size.
 Different Candidate Sizes.
 In the proposed two-stage framework SA-Person, the candidate set produced by the appearance-based retrieval stage serves as the input for the SceneRanker to perform scene-aware reranking.
-The size of this set ($K$) directly affects the trade-off between including the correct match and introducing noisy distractors.
-To better understand this trade-off and determine an optimal value for ($K$), we conduct a parameter analysis to systematically
-Figure 3: Impact of candidate size.
-investigate the impact of different candidate set sizes on the final retrieval performance.
-As shown in Figure  3, the performance of SceneRanker exhibits a trend of first increasing and then decreasing as the candidate size changes.
+The size of this set ($K$) governs the trade-off between ensuring the inclusion of the ground-truth match and minimizing the introduction of noisy distractors.
+To better understand this trade-off and determine an optimal value for ($K$), we conduct a parameter analysis to systematically investigate the impact of different candidate set sizes on the final retrieval performance.
+As shown in Figure  6, the performance of SceneRanker exhibits a trend of first increasing and then decreasing as the candidate size changes.
 The best performance is achieved when the candidate size is set to 10.
@@
 These results indicate that a moderate candidate pool is optimal, as it balances retaining the ground-truth instance while minimizing noise, ultimately enhancing retrieval accuracy.
-4.5 Qualitative Results
-To further validate the effectiveness of SA-Person, we present a qualitative comparison in Figure 4.
-IRRA focuses solely on pedestrian appearance, retrieving candidates with similar coarse attributes, but fails to utilize scene information and thus struggles to distinguish individuals in different contexts.
-LongCLIP, on the other hand, leverages scene-level information to retrieve images consistent with the described environment, but lacks fine-grained attention to pedestrian appearance and often returns individuals who differ significantly from the target identity.
-In comparison, SA-Person first retrieves candidates based on pedestrian appearance and then employs an MLLM to understand scene context and re-rank these candidates accordingly.
-This approach enables SA-Person to more accurately align with both the appearance and context described in the query, leading to significantly improved retrieval accuracy over existing baselines.
-Figure 4: Qualitative comparison of retrieval results across three methods.
-The leftmost column shows the textual query and its corresponding ground-truth image.
-The top-6 retrieval results are presented from top to bottom for IRRA (ViT-L/14), LongCLIP (ViT-L/14), and SA-Person (ours).
-5 Conclusion
+VI-E Versatility Across Diverse Retrieval Scenarios
+To demonstrate the versatility of SA-Person, we evaluate its performance across three distinct retrieval scenarios: person retrieval in surveillance scenarios using augmented CUHK-SYSU from our preliminary analysis in Section III, person retrieval in varied real-world environments using our SCENEPERSON-13W benchmark, and general image-text retrieval in open-domain contexts using Urban1K [zhang2024long] and a 5K subset of ShareGPT4V [chen2024sharegpt4v].
+TABLE VII: SceneRanker Enhancements in Surveillance Scenarios.
+Methods R@1 R@5 R@10 mAP
+IRRA (ViT-B/16) [jiang2023cross] 47.60 84.00 89.14 63.63
++SceneRanker $54.24_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+6.64}}$ $86.35_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+2.35}}$ $89.14_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+0.00}}$ $68.03_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+4.40}}$
+IRRA (ViT-L/14) [jiang2023cross] 50.57 85.54 89.36 66.01
++SceneRanker $57.17_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+6.60}}$ $87.19_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+1.65}}$ $89.36_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+0.00}}$ $70.04_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+4.03}}$
+RDE (ViT-B/16) [qin2024noisy] 48.73 83.78 88.88 63.93
++SceneRanker $56.51_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+7.78}}$ $6.17_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+2.39}} 88$$8.88_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+0.00}}$ $69.27_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+5.34}}$
+TABLE VIII: SceneRanker Enhancements in General Image-Text Retrieval Scenarios.
+Model (Backbone) Urban1k shareGPT4V
+R@1 R@5 R@10 mAP R@1 R@5 R@10 mAP
+CLIP (ViT-B/16) 53.10 78.40 85.20 64.34 47.60 73.92 83.12 59.72
++ SceneRanker 72.70 83.00 85.20 77.69 59.52 78.40 83.10 68.19
+CLIP (ViT-L/14) 55.90 79.50 86.50 66.81 51.06 74.72 83.22 61.87
++ SceneRanker 73.10 83.90 86.50 78.36 61.34 78.94 83.22 69.53
+Figure 7: Qualitative Comparison of Retrieval Results with and without SceneRanker.
+The leftmost column shows the text and ground-truth image.
+The green frame contains the retrieval results from the appearance-based stage before SceneRanker re-ranking, while the blue frame contains the retrieval results after SceneRanker re-ranking.
+Figure 8: Loglog Plot of Inference Time Across Gallery Sizes. The plot shows inference time (normalized units, log scale) versus gallery size NN (log scale) for three methods.
+Table VII presents SceneRanker enhancements on augmented CUHK-SYSU in surveillance scenarios.
+It achieves R@1 gains ranging from 6.60% to 7.78% and mAP improvements from 4.03% to 5.34% on IRRA and RDE backbones, demonstrating the module’s robustness in complex surveillance environments.
+On SCENEPERSON-13W, SA-Person yields large improvements over baselines as shown in Tables IV.
+For Urban1K and ShareGPT4V, we integrate plug-and-play SceneRanker with CLIP baselines using full images and enriched descriptions.
+Table VIII reports retrieval performance on these datasets.
+SceneRanker enhances the baselines across both: R@1 improves by over 19% and mAP by more than 12% on Urban1K, effectively exploiting urban contextual cues such as landmarks and crowds.
+Comparable gains appear on ShareGPT4V, with $\sim$12% R@1 and $\sim$8% mAP uplifts, affirming adaptability to open-domain diversity.
+These results across scenarios confirm SA-Person’s broad applicability, extending beyond SCENEPERSON-13W to multimodal tasks requiring scene-aware disambiguation.
+VI-F Computational cost of inference
+Unlike prior approaches [niu2025chatreid] that rely on exhaustive MLLM-based retrieval across the entire gallery, SA-Person combines the efficiency of domain-specific models for appearance grounding with the robust contextual reasoning of MLLMs, achieving a strong balance between high retrieval performance and computational scalability.
+As defined in Section III, let $\mu_{s}$ and $\mu_{m}$ denote the per-image time units for domain-specific models and MLLMs, respectively, with $\mu_{m}\approx 2$ orders of magnitude larger than $\mu_{s}$ due to intensive multimodal processing.
+For a gallery of size $N$, Appearance-based Retrieval processes all $N$ full-scene images to yield $O(N\mu_{s})$ time overall, producing a compact set of $K$ candidates ($K=10\ll N$).
+Scene-aware Re-ranking then applies a single MLLM forward pass on these $K$ candidates with bounding boxes, incurring $O(K\mu_{m})$ time.
+The total time simplifies to $O(N\mu_{s}+K\mu_{m})$, orders of magnitude lower than gallery-wide MLLM retrieval at $O(N\mu_{m}+M\mu_{m})$ (where $K=10$, $M\ll N$), while surpassing pure domain-specific models at $O(N\mu_{s})$ through contextual re-ranking.
+As illustrated in Figure 8, this design renders the MLLM stage independent of $N$, ensuring scalability for large galleries.
+VI-G Qualitative Results
+To further validate the effectiveness of the proposed SA-Person framework, we present a qualitative comparison of retrieval results before and after applying the SceneRanker module in Figure 7.
+Without scene-aware re-ranking, the initial candidates retrieved by the appearance-based methods often include individuals with similar coarse attributes but in mismatched contexts, leading to suboptimal ranking.
+In contrast, SA-Person leverages an MLLM to interpret global scene semantics and re-rank these candidates accordingly.
+These qualitative examples demonstrate that SA-Person effectively bridges the gap between fine-grained appearance matching and holistic scene understanding, achieving more accurate and robust text-based person retrieval.
+VII Limitations and Discussion
+While SA-Person advances scene-aware text-based person retrieval, our approach has several limitations.
+First, the textual descriptions in SCENEPERSON-13W are generated via MLLMs to enable large-scale annotation, which efficiently captures diverse appearance and contextual details but may exhibit a relatively uniform stylistic phrasing compared to the diverse variations in real-world texts.
+A more effective strategy may involve incorporating diverse templates during generation to broaden stylistic coverage, and exploring directly optimizing the MLLM to select the most varied phrasing dynamically.
+Additionally, SceneRanker builds on off-the-shelf MLLMs in a training-free manner for plug-and-play flexibility and broad applicability, yet task-specific fine-tuning could amplify its precision in retrieval intent alignment and scene context disambiguation.
+Exploring lightweight adapters or domain-adaptive pretraining would offer a promising avenue to boost performance without sacrificing zero-shot robustness.
+These directions not only address potential refinements but also open avenues for extending SA-Person to interactive human-centric systems and dynamic video retrieval.
+VIII Conclusion
 In this paper, we tackle the challenging task of person retrieval by leveraging multimodal information from both pedestrian appearance and global scene context.
-Specifically, we make two key contributions.
-First, we present ScenePerson-13W, the first large-scale dataset to offer comprehensive multimodal annotations, integrating fine-grained pedestrian appearance descriptions with rich scene-level contextual information within full-scene images.
-Second, we propose SceneRanker, a novel scene-aware re-ranking method that exploits multimodal cues from the entire scene to effectively refine retrieval results.
-Building upon these contributions, we develop SA-Person, a two-stage framework that initially grounds discriminative appearance features in cropped pedestrian regions and subsequently applies SceneRanker, a training-free multimodal large language model—to perform scene-aware ranking based on holistic image context.
+Specifically, we make three key contributions.
+First, we present ScenePerson-13W, the first large-scale dataset offering comprehensive multimodal annotations that integrate fine-grained pedestrian appearance descriptions with rich scene-level contextual information within full-scene images.
+Second, we propose SA-Person, a two-stage retrieval framework that combines discriminative appearance grounding in cropped pedestrian regions with holistic scene understanding for robust retrieval from large-scale full-scene databases.
 Extensive experimental results demonstrate that SA-Person significantly outperforms existing baselines, underscoring the critical importance of integrating both local appearance cues and global scene context for robust and effective text-based person retrieval in complex real-world scenarios.
-Limitations.
-While our work advances scene-aware person retrieval, it has two main limitations.
-First, the query descriptions in ScenePerson-13W are generated by MLLM, which may slightly deviate from natural human expression in retrieval scenarios.
-Second, our SceneRanker builds on off-the-shelf MLLMs without task-specific training, leaving room for improvement through fine-tuning to enhance alignment with retrieval intent and strengthen scene context understanding.
-Appendix A Appendix of SA-Person: Text-Based Person Retrieval with Scene-aware Re-ranking
-In the appendix, we provide the following materials:
-  • Section A.1 provides a discussion of related work on text-based person retrieval and multimodal large language models.
-  • Section A.2 analyzes the broader impact of scene-aware person retrieval, addressing societal benefits and ethical considerations.
-  • Section A.3 details the construction and description generation processes for the SCENEPERSON-13W dataset.
-  • Section A.4 presents supplemental results based on varisou retrieval methods and prompt designs, and visualizations of SceneRanker’s performance in Section 4.4 and 4.5 in the main paper.
-  • Section A.5 describes prompt designs and instructional templates for SceneRanker in Section 4.4 in the main paper.
-A.1 Related works
-Text-Based Person Retrieval.
-Text-Based Person Retrieval (TBPR) aims to retrieve specific individuals based on natural language descriptions, typically from a gallery of cropped person images [45, 19, 55, 26, 25].
-These methods focus on fine-grained cross-modal alignment between visual appearances and textual attributes, such as clothing, gender, or accessories [34, 55, 37, 38, 51].
-With the advent of vision-language models like CLIP [31], TBPR performance has significantly improved.
-For instance, IRRA [15] introduces a multi-modal interaction encoder for implicit relational reasoning, and TBPS-CLIP [5] proposes a simple yet strong baseline through enhanced training strategies.
-To address noise in cross-modal data, RaSa [3] proposes sensitivity-aware and relation-aware representation learning for robustness against weakly matched pairs; MARS [10] incorporates visual reconstruction and attribute-sensitive losses to mitigate inter-identity noise and intra-identity variations; RDE [30] introduces dual embeddings to handle under-correlated or mislabeled image-text pairs under noisy conditions.
-Beyond cropped-image retrieval, some methods [50, 35, 42] retrieve individuals directly from unprocessed full-scene images, where individuals need to be located and identified in complex visual environments.
-Specifically, Zhang et al.
-[50] integrates textual cues into region proposals, enabling text descriptions to guide the detection process.
-Su et al.
-[35] introduces a memory-aided coarse-to-fine alignment strategy, using memory banks to enhance cross-modal contrastive learning.
-Yan et al.
-[42] explores multimodal fusion by combining textual and visual features with precise segmentation techniques to boost retrieval performance.
-While these methods often overlook the rich contextual information present in the surrounding scene, which could offer valuable clues for more accurate and robust person retrieval.
-Our SA-Person framework distinguishes itself by explicitly integrating scene context with appearance-based retrieval, leveraging textual descriptions of both individuals and their surroundings to enhance accuracy and disambiguate similar-looking candidates.
-Multimodal Large Language Models for Scene and Region-Level Reasoning.
-Recent advances in Multimodal Large Language Models (MLLMs) enabled powerful visual reasoning across complex scenes.
-Current efforts primarily progresses along two directions: comprehensive scene understanding using global image-text inputs [7, 8, 2, 36, 28, 11], and fine-grained region-level grounding [13, 54, 29, 6, 27, 32].
-For comprehensive scene understanding, recent works [28, 11] enable MLLMs to explicitly represent structured and compositional semantics.
-For instance, Mitra et al.
-[28] proposes generating scene graphs via MLLMs and incorporating them into prompts, facilitating compositional reasoning without relying on annotated scene graphs; MLLM-SUL [11] introduces a dual-branch encoder and a fine-tuned LLaMA [36] model to jointly perform scene description in traffic scenarios.
-To support region-level grounding, recent studies [13, 54, 6] enhance MLLMs with explicit spatial conditioning.
-For example, RegionGPT [13] and RegionBLIP [54] introduce region-wise feature pooling within bounding boxes or masks to extract localized representations for improved spatial reasoning; KOSMOS-2 [29] and Shikra [6] propose to encode spatial positions as discrete tokens or numeric coordinates to support referential understanding; GLaMM [32] facilitates grounded interaction through joint generation of text and segmentation masks.
-In this work, we explore the application of MLLMs to fine-grained person retrieval in complex scenes, where both textual descriptions and full-scene images are jointly leveraged to assess and rank visual relevance.
-A.2 Broader impacts
-The SA-Person framework and SCENEPERSON-13W dataset advance text-based person retrieval, improving public safety, smart city accessibility, and multimodal AI research.
-However, their surveillance applications risk privacy violations and biases from YouTube-derived data.
-To mitigate, restricting access with ethical guidelines ensures only responsible, research-focused use, preventing misuse by malicious actors and aligning with legal standards.
-Enhancing robustness through fine-tuning improves model accuracy across diverse scenarios, reducing errors and increasing reliability, thus maximizing societal benefits while ensuring ethical deployment.
-Figure 5: Overview of person retrieval limitation and scene-aware insights.
-(a) Illustration of appearance-based retrieval limitation, where multiple candidates match the query “A girl wearing a green and white striped shirt is smiling” due to similar appearances.
-(b) Illustration of scene-aware reranking, ranking candidate images by aligning scene context like “located in a restaurant, near a window with red text” within the query.
-(c) Wordcloud of contextual elements in ScenePerson-13W, highlighting diverse scene descriptors.
-(d) Examples from the ScenePerson-13W dataset, showcasing detailed scene and relational descriptions for person retrieval.
-Figure 6: Example instances from the ScenePerson-13W dataset.
-Descriptions in green text detail pedestrian appearance, while those in blue text describe their interactions within complex scenes.
-To clearly identify the target pedestrian associated with each textual description in scenes containing multiple pedestrians, the target individual is highlighted with a red bounding box in these visualizations.
-A.3 More Details of ScenePerson-13W
-The ScenePerson-13W dataset addresses critical challenges in scene-aware person retrieval by integrating detailed pedestrian appearance with rich scene context, overcoming limitations of appearance-only methods, as illustrated in Figure 5.
-This section provides additional details on the construction pipeline of the SCENEPERSON-13W dataset, with a focus on specific methodologies and confidence thresholds to ensure reproducibility and clarity.
-Pedestrian Detection and Tracking. Using the YOLOv11x, we detect and track pedestrians in the QVHighlights video dataset [17] with a confidence threshold of 0.6 to balance precision and recall.
-Completeness Filtering.
-To ensure quality, incomplete detections are filtered using MediaPipe’s pose estimation [1] with a minimum detection confidence of 0.65.
-Head detection requires at least one keypoint (e.g., nose, eyes) with visibility over 0.4, and shoulders need visibility over 0.6 for at least one shoulder.
-Only images with a head and one shoulder are retained.
-Deduplication.
-Redundancy is reduced using CLIP [31] to compute cosine similarity, normalized to the range [0, 1], for cropped pedestrian regions and full-scene images.
-Images with similarity score exceeding 0.91 are deemed redundant, and one instance is kept, yielding 131,438 unique pedestrian instances for diversity.
-Testing Set Refinement.
-To create a fair and robust testing set of 10,000 images, we address redundancy from frame-based tracking, where pedestrians appear across consecutive frames with minor variations.
-We apply CLIP-based visual clustering using the clip-vit-base-patch32 [31] model, with a similarity threshold of 0.88.
-Images are clustered via agglomerative clustering with average linkage, and one representative image per cluster is selected.
-This deduplicated testing set balances diversity and manageability, preserving complex scene and identity distributions for meaningful retrieval evaluation.
-Description Generation.
-Textual descriptions for each of the 131,438 pedestrian instances are generated using InternVL2.5-8B [8].
-The model is prompted with a full-scene image, highlighting the target pedestrian with a red bounding box, using the following prompt:
-Generate a description of the individual within the red bounding box and the connection with the surroundings, in a style similar to the template: template.
-Ensure that the description is concise within 50 words and accurate.
-Do not include text details or labels, and strictly avoid any punctuation other than commas and periods.
-The template guides the model to describe: (1) the individual’s appearance, (2) their location in the scene, and (3) relationships with surroundings. Specifically, the template is defined as:
-A [man/woman/girl/boy] wearing [describe basic clothing, notable accessories, and physical attributes] is [describe their action or posture].
-This person is located [describe their location, highlighting any landmarks, structures, or recognizable surroundings].
-Next to/to the left of/to the right of/behind/in front of/under/above/around [the man/woman/etc.], there is/are [briefly describe nearby people, objects, or structures, such as another person in red holding something, a group of people talking, or a pedestrian bridge].
-This structured approach ensures descriptions capture both pedestrian details and contextual cues, as illustrated in Figure 6.
-Dataset Examples.
-More examples of ScenePerson-13W are shown in Figure 6.
-showcasing the dataset’s rich scene and relational descriptions.
-These examples highlight the dataset’s ability to provide fine-grained appearance details and diverse scene contexts, addressing challenges like those in Figure 5, where scene-aware insights disambiguate similar-looking individuals.
-A.4 Experiments and Visualisation
-In this section, we provide deeper insights into our experimental results, ablation findings, and qualitative visualizations.
-A.4.1 SceneRanker Performance Across Retrieval Methods and Prompt Designs
-Table 6: Performance of SceneRanker with various text-based person retrieval methods under different prompt designs. Red numbers indicate the improvement brought by each prompt design.
-     Methods      R@1      R@5      R@10      mAP
-     TBPS-CLIP [5]      50.29      75.24      81.58      61.54
-      +NP      $71.41_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+21.12}}$      $80.19_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+4.95}}$      $81.58_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+0.00}}$      $76.03_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+14.49}}$
-      +BOP      $70.77_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+20.48}}$      $80.16_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+4.92}}$      $81.58_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+0.00}}$      $75.69_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+14.15}}$
-      +BEP      $\textbf{72.25}_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+21.96}}$      $\textbf{80.23}_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+4.99}}$      $\textbf{81.58}_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+0.00}}$      $\textbf{76.54}_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+15.00}}$
-     IRRA(ViT-B/16) [15]      60.57      83.35      87.86      70.65
-      +NP      $76.80_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+16.23}}$      $86.41_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+3.06}}$      $87.86_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+0.00}}$      $81.53_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+10.88}}$
-      +BOP      $75.78_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+15.21}}$      $86.39_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+3.04}}$      $87.86_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+0.00}}$      $80.99_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+10.34}}$
-      +BEP      $\textbf{77.42}_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+16.85}}$      $\textbf{86.61}_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+3.26}}$      $\textbf{87.86}_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+0.00}}$      $\textbf{81.96}_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+11.31}}$
-     IRRA(ViT-L/14) [15]      63.37      84.54      88.67      72.71
-      +NP      ${77.55}_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+14.18}}$      ${87.42}_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+2.88}}$      ${88.67}_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+0.00}}$      ${82.33}_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+9.62}}$
-      +BOP      ${76.69}_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+13.32}}$      ${87.29}_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+2.75}}$      ${88.67}_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+0.00}}$      ${81.85}_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+9.14}}$
-      +BEP      $\textbf{78.34}_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+14.97}}$      $\textbf{87.54}_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+3.00}}$      $\textbf{88.67}_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+0.00}}$      $\textbf{82.86}_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+10.15}}$
-     RDE [30]      61.26      83.50      88.23      71.17
-      +NP      $77.13_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+15.87}}$      $86.83_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+3.33}}$      $88.23_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+0.00}}$      $81.87_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+10.70}}$
-      +BOP      $76.29_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+15.03}}$      $86.88_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+3.38}}$      $88.23_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+0.00}}$      $81.46_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+10.29}}$
-      +BEP      $\textbf{77.58}_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+16.32}}$      $\textbf{86.84}_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+3.34}}$      $\textbf{88.23}_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+0.00}}$      $\textbf{82.18}_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+11.01}}$
-     RaSa [3]      62.51      83.83      88.09      72.05
-      +NP      $77.04_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+14.53}}$      $86.66_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+2.83}}$      $88.09_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+0.00}}$      $81.76_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+9.71}}$
-      +BOP      $76.24_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+13.73}}$      $86.74_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+2.91}}$      $88.09_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+0.00}}$      $81.29_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+9.24}}$
-      +BEP      $\textbf{77.84}_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+15.33}}$      $\textbf{86.83}_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+3.00}}$      $\textbf{88.09}_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+0.00}}$      $\textbf{82.27}_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+10.22}}$
-     MARS [10]      62.93      84.26      88.65      72.43
-      +NP      $77.26_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+14.33}}$      $87.30_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+3.04}}$      $88.65_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+0.00}}$      $82.10_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+9.67}}$
-      +BOP      $76.57_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+13.64}}$      $87.14_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+2.88}}$      $88.65_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+0.00}}$      $81.69_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+9.26}}$
-      +BEP      $\textbf{78.32}_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+15.39}}$      $\textbf{87.39}_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+3.13}}$      $\textbf{88.65}_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+0.00}}$      $\textbf{82.77}_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+10.34}}$
-To further assess SceneRanker’s generalizability, we evaluate its performance when integrated with various text-based person retrieval methods, each paired with three prompt designs: Naive Prompt (NP), Box Overlay Prompt (BOP), and Box Embedded Prompt (BEP).
-Results are presented in Table 6.
-SceneRanker significantly boosts performance for all methods, with R@1 improvements exceeding above 14.97% when using BEP, underscoring its compatibility and effectiveness in refining appearance-based retrieval.
-Across all methods, BEP consistently outperforms NP and BOP, achieving the highest R@1, suggesting that embedding bounding box coordinates enhances spatial grounding, enabling better alignment with textual descriptions.
-A.4.2 Visualization of Retrieval Results
-To illustrate the impact of SceneRanker, we present a qualitative comparison of retrieval results before and after applying SceneRanker, as shown in Figure 7.
-Figure 7: Qualitative Comparison of Retrieval Results with and without SceneRanker.
-The leftmost column shows the query and ground-truth image.
-The green frame contains the retrieval results from the appearance-based stage before SceneRanker re-ranking, while the blue frame contains the retrieval results after SceneRanker re-ranking.
-A.5 SceneRanker Prompt Design and Templates
-This subsection details the instructional prompt templates used in the SceneRanker ablation studies.
-SceneRanker performs reranking without task-specific fine-tuning, using a multimodal prompt comprising a visual prompt $\mathbf{P}_{v}$ and an instructional textual prompt.
-Within each prompt strategy, $\mathbf{P}_{v}$ is fixed, but the instructional prompt varies to enhance spatial grounding, with multiple prompts tested.
-We describe: (1) the instructional prompt templates for three strategies evaluated for enhancing spatial grounding in SceneRanker, each in a dedicated table, and (2) the model-specific instructional prompt templates designed for Qwen-VL 2.5 [2], MiniCPM-V-2_6 [46], and InternVL 2.5 [8] in the MLLMs ablation study, tailored to each model’s capabilities while ensuring fair comparison.
-All prompts are designed for use with the InternVL2.5-8B model [8] unless otherwise specified.
-A.5.1 Scene-aware Multimodal Prompt Templates
-For each prompt strategy, the visual prompt $\mathbf{P}_{v}$ is fixed, but multiple instructional prompts were tested to optimize spatial grounding of pedestrian-specific region information.
-Below, we specify $\mathbf{P}_{v}$ for each strategy, describe the instructional prompt variations for reproducibility, and present examples in dedicated tables, with the final adopted prompt in bold.
-[…]
-"For the text description: ’text’, order the images based on how accurately they reflect overall context in the text, starting with the most faithful match.
-Please output only a Python list format like this: [4, 2, 8, 1, 5, 3, 6, 9, 10, 7]." "For the description: ‘$T$’, 10 images are provided.
-Each image includes one person who is the intended candidate subject described in the text.
-Order the images based on how accurately they reflect overall context in the text, starting with the most faithful match.
-Please output only a Python list format like this: [4, 2, 8, 1, 5, 3, 6, 9, 10, 7]." "For the description: ‘$T$’, 10 images are provided.
-In each image, one person is implicitly assumed to be the potential match for the individual described.
-Order the images based on how accurately they reflect overall context in the text, starting with the most faithful match.
-Please output only a Python list format like this: [4, 2, 8, 1, 5, 3, 6, 9, 10, 7]." "For the given description: ‘$T$’, analyze the 10 provided candidate images.
-Each image contains one person implicitly corresponding to a potential subject referenced in the description.
-Evaluate how accurately each candidate aligns with the contextual details, visual cues, and semantic elements in the text description.
-Rank these images in descending order of relevance and contextual fidelity, with the most accurate match first.
-Return only a Python list of indices in the format: [4, 2, 8, 1, 5, 3, 6, 9, 10, 7], where each number represents the original position of the image in the candidate set." "Given the description: ‘$T$’, 10 images are provided, each containing a person implicitly regarded as the candidate subject described in the text.
-Reorder the images based on how well they match the description, considering both the assumed subject’s appearance and the surrounding context.
-Provide the output as a Python list in descending order of relevance, formatted like this: [4, 2, 8, 1, 5, 3, 6, 9, 10, 7]."
-Table 7: Naive Prompt (NP) instructional prompt templates for SceneRanker.
-"For the text description: ‘$T$’, order the images based on how accurately they reflect overall context in the text, starting with the most faithful match.
-Please output only a Python list format like this: [4, 2, 8, 1, 5, 3, 6, 9, 10, 7]." "For the description: ‘$T$’, 10 images are provided.
-Each image includes one person highlighted with a bounding box, representing the candidate subject mentioned in the text.
-Order the images based on how accurately they reflect overall context in the text, starting with the most faithful match.
-Please output only a Python list format like this: [4, 2, 8, 1, 5, 3, 6, 9, 10, 7]." "For the description: ‘$T$’, 10 images are provided.
-In each image, a person is marked with a bounding box as the potential match for the individual described.
-Order the images based on how accurately they reflect overall context in the text, starting with the most faithful match.
-Please output only a Python list format like this: [4, 2, 8, 1, 5, 3, 6, 9, 10, 7]." "For the given description: ‘$T$’, analyze the 10 provided candidate images.
-Each image contains one person highlighted with a bounding box, representing a potential subject referenced in the description.
-Evaluate how accurately each candidate aligns with the contextual details, visual cues, and semantic elements in the text description.
-Rank these images in descending order of relevance and contextual fidelity, with the most accurate match first.
-Return only a Python list of indices in the format: [4, 2, 8, 1, 5, 3, 6, 9, 10, 7], where each number represents the original position of the image in the candidate set." "Given the description: ‘$T$’, 10 images are provided, each containing a person highlighted with a bounding box, representing the candidate subject described in the text.
-Reorder the images based on how well they match the description, considering both the highlighted subject’s appearance and the surrounding context.
-Provide the output as a Python list in descending order of relevance, formatted like this: [4, 2, 8, 1, 5, 3, 6, 9, 10, 7]."
-Table 8: Box Overlay Prompt (BOP) instructional prompt templates for SceneRanker.
-"For the text description: ‘$T$’, order the images based on how accurately they reflect overall context in the text, starting with the most faithful match.
-Please output only a Python list format like this: [4, 2, 8, 1, 5, 3, 6, 9, 10, 7]." "For the text description: ‘$T$’, 10 images are provided.
-Each image includes one person marked by a bounding box (given as coordinates in <box>).
-This bounding box indicates the specific person selected as a potential match for the description.
-Please rank the images based on how accurately the boxed person and their surrounding scene match the description, starting with the most accurate match.
-Please output only a Python list format like this: [4, 2, 8, 1, 5, 3, 6, 9, 10, 7]." "For the text description: ‘$T$’, 10 images are provided.
-Each image includes a bounding box, which indicates the specific person selected as a potential match for the description, and its coordinates are shown as <box> for each image.
-Please rank the images based on how accurately the boxed person and their surrounding scene match the description, starting with the most accurate match.
-Please output only a Python list format like this: [4, 2, 8, 1, 5, 3, 6, 9, 10, 7]." "For the description: ‘$T$’, 10 images are provided.
-In each image, one person is implicitly assumed to be the potential match for the individual described.
-Order the images based on how accurately they reflect overall context in the text, starting with the most faithful match.
-Please output only a Python list format like this: [4, 2, 8, 1, 5, 3, 6, 9, 10, 7]." "For the description: ‘$T$’, 10 images are provided.
-Each image includes one person who is the intended candidate subject described in the text.
-Order the images based on how accurately they reflect overall context in the text, starting with the most faithful match.
-Please output only a Python list format like this: [4, 2, 8, 1, 5, 3, 6, 9, 10, 7]."
-Table 9: Box Embedded Prompt (BEP) instructional prompt templates for SceneRanker.
-[…]
-[…]
-A.5.2 Model-Specific Instructional Prompt Templates
-For the MLLM ablation study, SceneRanker is evaluated with Qwen-VL 2.5, MiniCPM-V-2_6, and InternVL 2.5, each using a tailored instructional prompt template paired with plain full-scene images.
-These prompts are designed to leverage each model’s strengths while ensuring fair comparison.
-Their respective instructional prompt templates are provided in Table 10.
-Qwen-VL 2.5: "For the text description: ‘$T$’, analyze the 10 candidate images and rank them from most to least consistent with the text.
-Output only a Python list containing all integers from 1 to 10, with no duplicates or omissions, like [4, 2, 8, 1, 5, 3, 6, 9, 10, 7]." MiniCPM-V: "Considering the full context of the description ‘$T$’, rank the provided images from best to worst match.
-Output only a Python list containing numbers from 1 to 10." InternVL 2.5: "For the text description: ‘$T$’, order the images based on how accurately they reflect overall context in the text, starting with the most faithful match.
-Please output only a Python list format like this: [4, 2, 8, 1, 5, 3, 6, 9, 10, 7]."
-Table 10: Instructional Prompt Templates for Comparative Evaluation of Multimodal Large Language Models.
+Acknowledgments
+This work was supported in part by the InnoHK Program of the Hong Kong SAR Government; the National Natural Science Foundation of China under Grants No.
+62306313, 62476188 and 62206276; the Natural Science Foundation of the Jiangsu Higher Education Institutions of China; and the Key Laboratory of New Generation Artificial Intelligence Technology & Its Interdisciplinary Applications (Southeast University), Ministry of Education, China.